### v1

# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

Abstract
Transformer-based models have significantly advanced time series forecasting, with patch-based input strategies offering efficiency and improved long-horizon modeling. Yet, existing approaches rely on temporally-agnostic patch construction, where arbitrary starting positions and fixed lengths fracture temporal coherence by splitting natural transitions across boundaries. This naive segmentation often disrupts short-term dependencies and weakens representation learning. In response, we propose EntroPE (Entropy-Guided Dynamic Patch Encoder), a novel, temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries. This preserves temporal structure while retaining the computational benefits of patching. EntroPE consists of two key modules, namely an Entropy-based Dynamic Patcher (EDP) that applies information-theoretic criteria to locate natural temporal shifts and determine patch boundaries, and an Adaptive Patch Encoder (APE) that employs pooling and cross-attention to capture intra-patch dependencies and produce fixed-size latent representations. These embeddings are then processed by a global transformer to model inter-patch dynamics. Experiments across long-term forecasting benchmarks demonstrate that EntroPE improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling. Code is available at https://github.com/Sachithx/EntroPE.
1 Introduction
Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer & Krauss, 2018), healthcare monitoring (Reis & Mandl, 2003), and environmental surveillance (Smith, 1989). However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim & Zohren, 2021; Torres et al., 2021). The transformer architecture (Vaswani et al., 2017) has revolutionized sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series forecasting. Despite their promise in learning long-term temporal relationships, time series transformers remain suboptimal due to unique complexities such as non-stationarity, seasonality, varying temporal granularities (Wen et al., 2023; Zeng et al., 2023).
Recent advances in time series transformers include Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), which address long-sequence modeling challenges through sparse attention mechanisms, decomposition techniques, and improved positional encodings while operating on point-wise inputs. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and long-term forecasting performance. This demonstrated the critical impact of input tokenization strategies on transformer performance in time series. However, subsequent approaches have predominantly adopted similar temporally-agnostic patching schemes, which may not fully capture the temporal coherence and statistical properties inherent in time series data.
Such temporally-agnostic patching introduces critical technical challenges. Fixed-length patching creates inconsistent input representations between training and inference phases. During training, time series samples from different time periods result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another). However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift where the model encounters patch configurations it rarely or never saw during training. Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for underlying dynamics (see Fig. 1(a)). For example, a gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition and forecasting.
Figure 1: (a) Temporally-agnostic, and (b) Temporally-informed (proposed) patching with calculated entropies for ETTh1 sample.
These challenges directly impact model performance through two mechanisms. Fragmented temporal patterns lead to incomplete intra-patch representations, where semantically related time points are separated across different patches and processed independently. Additionally, the train-inference mismatch reduces the model’s ability to generalize, as it must extrapolate to patch configurations with different statistical properties than those encountered during training. This is particularly detrimental for capturing rapid transitions and fine-grained temporal dynamics that require consistent temporal context.
Unlike prior work that treats patching as a static preprocessing step, we propose EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates dynamic boundary detection as a core architectural component. Our approach employs entropy-based criteria to identify natural temporal transition points, ensuring patch boundaries align with the underlying temporal structure rather than arbitrary positions (see Figure 1(b)). We then employ adaptive encoding mechanisms to process these variable-length patches while preserving intra-patch dependencies. This temporally-informed approach addresses both the train-inference mismatch and boundary fragmentation issues inherent in temporally-agnostic patching methods.
The main contributions of this paper are summarized as follows:
  • We introduce the first information-theoretic approach to patching for time series forecasting, where conditional entropy is used to place boundaries that respect temporal causality and predictive difficulty.
  • We design an adaptive encoder that converts variable-length patches into fixed representations via pooling and cross-attention, enabling efficient batch computation.
  • We demonstrate consistent gains across diverse benchmarks, showing that entropy-guided dynamic patching is a practical and generalizable direction for time series transformers.
2 Related Work
Time Series Transformers and Long-Term Forecasting. The transformer architecture has fundamentally transformed sequence modeling across domains, with time series forecasting being no exception. Early adaptations like Informer (Zhou et al., 2021) addressed the quadratic complexity of self-attention in long sequences through ProbSparse attention mechanisms, while Autoformer (Wu et al., 2021) and FEDformer (Zhou et al., 2022) introduced decomposition-based approaches and frequency domain modeling to capture complex temporal dependencies. These foundational works established transformers as viable architectures for time series tasks but operated on point-wise inputs, leading to computational challenges and a limited ability to capture local temporal patterns efficiently. The fundamental challenge these approaches faced was the trade-off between sequence length and computational tractability. Point-wise processing, while preserving fine-grained temporal information, resulted in prohibitively long input sequences for practical applications, particularly in long-term forecasting scenarios where historical contexts spanning hundreds or thousands of time steps are crucial for accurate predictions.
Patch-Based Input Strategies. Inspired by vision transformers (Dosovitskiy et al., 2021), patch-based input strategies have been widely adopted in time series forecasting. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) segment time series into fixed-length patches that serve as input tokens. This design achieves a quadratic reduction in sequence length while preserving local semantic information through channel-independent processing, demonstrating that tokenization strategies can significantly improve both computational efficiency and forecasting accuracy. Building on this foundation, Crossformer (Zhang & Yan, 2023) introduced hierarchical mechanisms tailored for modeling both intra-patch and inter-patch dependencies. CARD (Xue et al., 2024) extended patching to multivariate forecasting, while patch-independence approaches (Lee et al., 2024) revealed superior representations in self-supervised settings which further highlights the importance of patch-based input strategies for time series modeling. Beyond Transformer architectures, xPatch (Stitsyuk & Choi, 2025) combined dual-flow MLP–CNN designs with exponential decomposition, whereas Pathformer (Chen et al., 2024) and PatchMLP (Tang & Zhang, 2025) leveraged hierarchical processing to capture multi-scale relationships.
However, fixed-length patching suffers from important limitations. Uniform segmentation can fragment coherent patterns such as seasonal transitions or trend shifts. During training, windows are sampled by shifting step by step so all possible boundaries are eventually seen, whereas inference fixes a single segmentation that may not align with natural transitions, weakening representation learning and harming generalization. In response, MSPatch (Cao et al., 2025) introduces multi-scale patch segmentation to capture both short- and long-term dependencies, mitigating the fragmentation of coherent temporal patterns. HDMixer (Huang et al., 2024) proposes a Length-Extendable Patcher (LEP) module that adjusts patch lengths via bi-linear interpolation to enrich boundary information, though it remains constrained by predetermined patch counts and may introduce interpolation-induced uncertainty. Beyond patch resizing, foundation models have explored alternative input constructions such as MOIRAI (Woo et al., 2024) leverages frequency-based patch sizing to preserve periodic structures, while Chronos (Ansari et al., 2024) bypasses patching entirely by treating individual time points as tokens, albeit at a substantially higher computational cost.
While recent works have advanced patch construction strategies, significant gaps remain. Existing adaptive methods often overlook causal structure and temporal coherence. This prevents them from capturing optimal boundaries guided by temporal uncertainty and predictive difficulty. Given the highly dynamic nature of time series across domains, a principled and fully adaptive approach is required. Moreover, efficient training capabilities are required, particularly due to the variable number of patches in each sample. In contrast, dynamic patching schemes in other domains, such as natural language processing, have shown superior performance and efficiency. For instance, the Byte Latent Transformer (Pagnoni et al., 2025) employes byte level inputs instead of traditional word tokenization, and dynamically decide patches as a part of the architeture. Motivated by these advances, our work introduces entropy-guided dynamic boundary detection that respects temporal causality and predictive uncertainty, coupled with adaptive encoding mechanisms specifically designed to preserve temporal dependencies within variable-length patches.
3 Methodology
3.1 Problem Formulation
Given a multivariate time series $X=[x_{1},x_{2},\ldots,x_{C}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, the goal is to predict $\hat{Y}\in\mathbb{R}^{C\times T}$ over forecast horizon $T$. Following the channel-independence principle from PatchTST (Nie et al., 2023), we deliberately disregard cross-channel dependencies and apply a shared function $f_{\phi}:\mathbb{R}^{L}\to\mathbb{R}^{T}$ (with shared parameters $\phi$) to each channel independently: $\hat{\mathbf{y}}_{c}=f_{\phi}(x_{c}),\quad c=1,2,\ldots,C$.
This design choice isolates the temporal modeling capabilities from cross-variate relationships, allowing us to focus purely on the effectiveness of our dynamic patching mechanism within each univariate series. The multivariate prediction is constructed by stacking channel-wise results:
$\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{C}]\in\mathbb{R}^{C\times T}$
Figure 2: Comprehensive architecture of EntroPE. The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate forecasting.
3.2 Forward Process
The forward process of EntroPE (see Fig. 2) transforms raw multivariate time series into forecasts through a sequence of interconnected components. The raw input multivariate time series $X\in\mathbb{R}^{B\times C\times L}$, where $B$, $C$, and $L$ represent batch size, number of channels (variables), and sequence length respectively, is first processed following the channel independence principle, converting it to $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$ for independent processing of each variable. As the next step, to mitigate the statistical difference between training and testing distributions, we integrate the Reversible Instance Normalization layer (Kim et al., 2021) which outputs $X^{Norm.}\in\mathbb{R}^{(B\cdot C)\times L}$.
Following the Chronos tokenization approach (Ansari et al., 2024) (Appendix A.6), we discretize the continuous input sequence into uniform bins based on a fixed-size vocabulary $V$, producing $X^{tok}$. The tokenized sequence $X^{tok}$ is then processed by the EDP (explained in Sec. 3.3), which generates a patch mask $M\in\mathbb{R}^{B\cdot C\times L}$ indicating optimal patch boundary locations. Using this patch mask, the original continuous input $X^{ci}$ is segmented into variable-length patches. These patches are fed into the APE (explained in Sec. 3.4), which generates patch embeddings while aggregating intra-patch dependencies.
The encoder produces two key outputs, namely, (1) patch embeddings $P^{in}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$, which serve as input embeddings for the global transformer, and (2) encoder hidden states $H^{enc}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ that capture temporal information at the original sequence resolution for subsequent decoding head. Here, $\hat{p}$ is not fixed due to the dynamic nature of the patching, $d_{t}$ and $d_{p}$ are time point-wise embedding and patch embedding dimensions, respectively.
The patch embeddings $P^{in}$ are processed by the Global Transformer, which learns inter-patch long-term dependencies and outputs enriched patch embeddings $P^{out}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$. Finally, the Fusion Decoder (FD) combines the global transformer outputs ($P^{out}$) with the encoder hidden states ($H^{enc}$) through cross-attention, flattens the enriched representations, and projects them through a linear head and Instance De-Normalization layer to generate predictions for the future $T$ time steps, producing the final forecast $\hat{Y}\in\mathbb{R}^{B\times C\times T}$.
3.3 Entropy-Based Dynamic Patcher (EDP)
As shown in Fig. 2(A), our dynamic patch boundary detection leverages a lightweight causal transformer to identify temporal transitions based on predictive uncertainty. The process operates in two independent phases, namely model pre-training for temporal pattern learning and entropy-guided patch boundary detection with frozen weights.
Entropy Model Pre-training.
Drawing inspiration from the decoder-only transformer architecture and cross-entropy training demonstrated in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens. Our lightweight architecture comprises 8 embedding dimensions (shared input and head), 2 transformer layers, 4 attention heads, totaling 3,648 learnable parameters. The model performs autoregressive prediction at time $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence at time $\{t_{1},t_{2},\ldots,t_{L-1}\}$ using standard cross-entropy loss.
This streamlined design efficiently captures temporal dynamics while maintaining computational tractability for entropy estimation. Crucially, our objective is not to achieve superior forecasting performance, but rather to align the causal transformer for measuring predictive uncertainty across the discrete vocabulary. This architectural approach enables identification of natural transition points in raw time series sequences, where predictive uncertainty that quantified through entropy over the output next-token distribution reaches local maxima.
Entropy-Guided Patch Boundary Detection. Upon training convergence, we freeze model weights and compute entropy for each time position. For position $t$, we calculate the Shannon entropy of the next-token prediction distribution:
$H(x_{t})=-\sum_{v\in\mathcal{V}}p_{\theta}(x_{t+1}=v|x_{\leq t})\log p_{\theta}(x_{t+1}=v|x_{\leq t})$ (1)
Following Pagnoni et al. (2025), we implement dual-threshold boundary detection combining global and relative uncertainty measures. A patch boundary is placed at position $t$ when three conditions are met simultaneously:
$H(x_{t})$ $>\theta\quad\text{(global entropy threshold)}$ (2)
$H(x_{t})-H(x_{t-1})$ $>\gamma\quad\text{(relative entropy increase)}$ (3)
This mechanism ensures patches begin only at positions with both high absolute uncertainty and significant uncertainty increases, avoiding fragmentation from minor entropy fluctuations while capturing meaningful temporal transitions. EDP outputs patch mask $M$, which indicates the non-overlapping patch segmentation for further processing.
3.4 Adaptive Patch Encoder (APE)
The EDP produces variable-length patches where both patch count $\hat{p}$ and individual lengths vary across samples (patch mask $M$). As shown in Fig. 2(B), the APE converts these heterogeneous patches into fixed-size representations suitable for transformer processing while preserving temporal information. When handling variable patches, we ensure batch processing compatibility by padding sequences to the maximum patch count within each batch and applying attention masks to handle variable patch lengths during cross-attention operations.
Architecture Design. Our encoder employs a two-stage approach. First initial dimensionality reduction via pooling (from time-point embeddings to initial patch embeddings), followed by iterative cross-attention refinement. This design is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), but specialized for temporal patch aggregation.
Initial Patch Embedding. We first obtain the time-point embedding $h_{0}$ for the entire sequence with length $L$ through an Embedding Layer $E\in\mathbb{R}^{V\times d_{t}}$ over the same Vocabulary and Tokenizer used in the Entropy Model. For each variable-length patch $p_{j}$, we generate initial patch representations via max pooling: $P_{0,j}=\text{MaxPool}(E_{f}(p_{j}))$, where $E_{f}$ represents the time-point embedding function. Max pooling captures salient temporal features while providing translation invariance within patches.
Cross-Attention Refinement. We refine patch representations using cross-attention, where patch embeddings serve as queries and time-point embeddings act as keys and values. For layer $n$:
$P_{n}=P_{n-1}+W_{o}\left(\text{softmax}\!\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right),$ (4)
with attention components:
$Q_{j}$ $=W_{q}(P_{n-1,j})\quad\text{(patch embeddings as queries)},$ (5)
$K_{i}$ $=W_{k}(h_{n-1,i}),\quad V_{i}=W_{v}(h_{n-1,i})\quad\text{(time-point embeddings as keys/values)},$ (6)
where $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ are learnable projection matrices corresponding to queries, keys, values, and output transformations, respectively. For $N$-layer architectures with $N>1$, time-point embeddings are iteratively refined through stacked transformer layers as $h_{n}=\text{APETransformer}(h_{n-1})$, with cross-attention mechanisms operating between consecutive layers. Here, $h_{n}$ is the time point embedding sequence at layer $n$.
To ensure that self-attention within each APETransformer operates exclusively within patch boundaries, we employ a tailored attention mask that accommodates the variable patch lengths throughout the sequence. This masking strategy prevents information leakage across patch boundaries while preserving intra-patch temporal dependencies.
The encoder outputs fixed-size patch embeddings $P^{in}=P_{N}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$ for global transformer processing and preserves encoder hidden states $H^{enc}=h_{N}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ for temporal information in the decoding stage.
3.5 Global Transformer & Fusion Decoder
The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate forecasts.
Global Transformer. The patch embeddings $P^{in}$ from the adaptive encoder are processed through a standard transformer architecture to capture long-range temporal relationships between patches. Since intra-patch dependencies are already modeled by the encoder, the global transformer focuses exclusively on inter-patch interactions. We employ bidirectional attention (non-causal) to enable comprehensive context modeling across the entire sequence: $P^{out}=\text{GlobalTransformer}(P^{in})$.
Fusion Decoder (FD). A key challenge in dynamic patching is that the number of patches $\hat{p}$ varies across samples, making direct forecasting from flattened patch representations problematic. To address this, we employ a cross-attention mechanism that fuses global patch context with fine-grained temporal information preserved from the encoder (see Fig. 2(C)). The decoder reverses the attention configuration used in the adaptive encoder. Here, encoder hidden states $H^{enc}$ serve as queries, while global patch embeddings act as keys and values:
$Q_{i}$ $=W_{q}(H^{enc}_{i})\quad\text{(time-point queries)},$ (7)
$K_{j}$ $=W_{k}(P^{out}_{j}),\quad V_{j}=W_{v}(P^{out}_{j})\quad\text{(patch keys/values)}$ (8)
The cross-attention operation enriches time-point representations with global context:
$H^{dec}=H^{enc}+W_{o}\left(\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right).$ (9)
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, ensuring preservation of both local temporal patterns and global contextual dependencies. Optionally, similar to APE, the iterative cross-attention refinement can be configured through alternating transformer layers and cross-attention layers, creating a structured information exchange between different representational levels.
[…]
4 Experiments
Datasets. We evaluate EntroPE on seven widely-used long-term multivariate forecasting benchmarks: ETT family (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, and Exchange Rate datasets, covering diverse domains and temporal characteristics (detailed descriptions in Appendix Tab. 4).
Baselines. We benchmark against 14 state-of-the-art deep forecasting models from 2021–2025, organized by architecture. (1) Transformer-based models: iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Kernel-attention model: TimeKAN (Huang et al., 2025b); (5) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (6) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025); (7) Filter-based model: FilterTS (Wang et al., 2025). (Further details on baselines are discussed in Appendix A.1.)
Experimental Settings. Following standard practices in time series forecasting literature, we evaluate performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics across multiple prediction horizons $T=\{96,192,336,720\}$ and fixed look-back window $L=96$. Adopting the same experiemnt setup as Huang et al. (2025b), baseline results up to 2024 are taken from TimeKAN, while other methods are re-evaluated using their official implementations. EntroPE experiemnts are conductd in 5 random seeds and we report the best results. (see Appendix A.3.)
4.1 Main Results
Table 1 summarizes average performance across all prediction horizons. EntroPE matches or surpasses, in many cases, recent variable-mixing, Transformer-based, CNN-based, MLP-based, foundation, lightweight, LLM-aligned, filter-based, and KAN-based models under the standard 96-step look-back with horizons up to 720. It retains PatchTST’s overall architecture but introduces two key innovations namely, (i) entropy-driven dynamic patching and (ii) adaptive embeddings. Relative to PatchTST, EntroPE achieves notable accuracy gains of approximately 20% on ETTh1, 15% on Electricity, and about 10% on average, while also reducing token count and improving efficiency by leveraging channel independence and non-overlapping patches. Beyond PatchTST, EntroPE consistently outperforms HDMixer’s length-extendable patching and improves upon TimeBase, though the latter remains highly efficient. On the Electricity dataset, iTransformer continues to offer the strongest performance-efficiency tradeoff given its channel-wise self-attention across 321 and 862 variables, respectively. CALF and LangTime also remain competitive, though their large model sizes (CALF has 18 million, LangTime has 500 million while EntroPE has 0.1 million range parameters) impose substantial costs. Across other benchmarks, EntroPE delivers robust improvements, underscoring the value of entropy-guided temporal segmentation for multivariate forecasting.
Table 1: Multivariate time series forecasting results on benchmark datasets. Results are averaged across prediction horizons $T=\{96,192,336,720\}$ with fixed input length $L=96$ for all datasets. Best results are highlighted in red and second-best in blue. A dash (-) indicates that the configuration was not found in the original implementation. (Full table: Appendix A.4)
Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338 0.613 0.539
FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327 0.519 0.429
DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319 0.354 0.414
TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304 0.416 0.443
PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318 0.367 0.404
Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299 0.338 0.391
HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295 - -
iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270 0.360 0.403
TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272 0.378 0.410
TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296 - -
LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285 0.361 0.400
TimeKAN [2025] 0.418 0.427 0.391 0.410 0.380 0.398 0.285 0.331 0.244 0.273 0.197 0.286 - -
FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275 0.355 0.400
CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266 - -
EntroPE [2025] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271 0.331 0.386
4.2 Ablation Studies
Model component analysis. To systematically evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720 time steps, measuring performance via MSE. We progressively remove components from our full EntroPE architecture to create four configurations: (1) EntroPE (Full): Dynamic patching + Adaptive encoder + Fusion decoder; (2) EntroPE - Dynamic: Static patching + Adaptive encoder + Fusion decoder; (3) EntroPE - (Dynamic Patching + Adaptive Encoder): Static patching + Max pooling + Fusion decoder; (4) EntroPE - (Dynamic Patching + Adaptive Encoder + Fusion Decoder): Static patching + Max pooling + Flattened output.
The full EntroPE architecture achieves the best performance across all datasets and horizons, while systematic component removal leads to progressive performance degradation (Table 2). The EntroPE - Dynamic Patching configuration yields second-best results, demonstrating the significant contribution of dynamic boundary detection. Further removing the adaptive encoder (EntroPE - (Dynamic Patching + Adaptive Encoder)) shows additional degradation, validating our cross-attention-based encoder design. The worst performance occurs when all three components are removed, producing only static patching with basic pooling and flattened output. This systematic performance decline confirms that each architectural choice addresses specific modeling challenges in time series forecasting.
Table 2: Ablation study showing component importance and dynamic vs. static patching comparison. Best results are highlighted in red and second-best in blue.
Dataset T Full -EDP -EDP -APE -EDP -APE -FD Baselines
Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat PatchTST iTrans.
ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519 0.501 0.487
720 0.439 0.460 0.469 0.477 0.461 0.527 0.503 0.500
ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462 0.427 0.428
720 0.397 0.448 0.470 0.463 0.466 0.456 0.436 0.427
ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427 0.421 0.426
720 0.445 0.451 0.459 0.452 0.457 0.483 0.462 0.491
Weather 336 0.258 0.261 0.265 0.270 0.262 0.262 0.284 0.278
720 0.341 0.343 0.353 0.359 0.347 0.346 0.356 0.358
Dynamic vs. Static Patching. In addition to the component ablations, we compare EntroPE with dynamic patching against three static (fixed-length, non-overlapping) patching schemes. Experiments are conducted on ETTh1, ETTm1, and Weather datasets with a fixed input length of 96 and forecasting horizons of 336 and 720. Dynamic patching achieves the lowest MSE in most settings (Table 2, while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme. However, the computational cost of this setting is substantially higher, as the self-attention mechanism must attend to every time step individually. Figure 3 illustrates the efficiency advantage of dynamic patching over static alternatives under different configurations.
Figure 3: Ablation study for Dynamic and Static(patch_length) patching with single epoch duration.
[…]
Figure 4: Ablation study for threshold sensitivity with epoch duration.
4.3 Efficiency Analysis
Figure 5: EntroPE efficiency analysis.
As shown in Figures 3 and 4, our model adapts to different patching schemes, directly affecting computational efficiency. In the dynamic setting, the model maintains superior predictive performance while operating within an optimal computation budget. We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), Mean Squared Error (MSE), and the number of trainable parameters. Figure 5 demonstrates that our model achieves a more favorable performance–efficiency trade-off, with bubble size proportional to the parameter count. Notably, while TimeMixer and TimeKAN occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture.
5 Conclusion
We introduced EntroPE, a temporally informed dynamic patch encoding framework that extends time series transformer architectures through entropy-guided boundary detection. Instead of partitioning sequences at arbitrary intervals, EntroPE strategically places patch boundaries at natural temporal transitions, preserving temporal coherence while supporting efficient variable-length patch processing. This design provides a principled way to incorporate information-theoretic signals into patching, yielding improved forecasting accuracy alongside practical computational benefits. Beyond its empirical performance, EntroPE highlights the importance of respecting the intrinsic structure of time series, opening new research directions in adaptive sequence modeling and context-aware temporal representation learning.
Appendix A Appendix
A.1 Baseline Details
Table 3: Overview of baseline models (2021–2025) grouped by architecture. We highlight their key design paradigms for deep time series forecasting.
Model Category Key Features / Innovations
iTransformer (Liu et al., 2024b) Transformer Inverted attention across feature and time dimensions.
PatchTST (Nie et al., 2023) Transformer Patch-based embedding with channel-independence.
FEDformer (Zhou et al., 2022) Transformer (freq.-decomp.) Fourier/Wavelet enhanced decomposition.
Autoformer (Wu et al., 2021) Transformer (decomp.-autocorr.) Progressive decomposition with auto-correlation.
TimesNet (Wu et al., 2023) CNN Temporal 2D variation modeling with convolutional layers.
TimeMixer (Wang et al., 2024) MLP Multiscale mixing, separation of trend and seasonal signals.
HDMixer (Huang et al., 2024) MLP Hierarchical dependency with extendable patches.
DLinear (Zeng et al., 2023) Linear/MLP Channel-independent trend/seasonal decomposition.
TimeKAN (Huang et al., 2025b) Kernel Attn. Network Kernelized attention distinct from Transformer/MLP.
Time-FFM (Liu et al., 2024a) Foundation Pre-trained foundation model for time series.
TimeBase (Huang et al., 2025a) Lightweight Resource-efficient, practical forecasting baseline.
LangTime (Niu et al., 2025) LLM-aligned / Cross-modal Language-guided forecasting with RL optimization.
CALF (Liu et al., 2025) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
FilterTS (Wang et al., 2025) Filter-based Static global + dynamic cross-variable frequency filtering.
A.2 Dataset Details
Table 4: Summary of datasets used for long-term forecasting evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
Dataset Dim Dataset Size (Train, Val, Test) Frequency Domain Information
ETTh1 7 (8545, 2881, 2881) Hourly Energy Infrastructure
ETTh2 7 (8545, 2881, 2881) Hourly Energy Infrastructure
ETTm1 7 (34465, 11521, 11521) 15min Energy Infrastructure
ETTm2 7 (34465, 11521, 11521) 15min Energy Infrastructure
Weather 21 (36792, 5271, 10540) 10min Meteorological Monitoring
Electricity (ECL) 321 (18317, 2633, 5261) Hourly Electricity Consumption
Exchange Rate 8 (5120, 665, 1422) Daily Foreign Exchange Market
This study employs a comprehensive collection of benchmark forecasting datasets to rigorously evaluate model performance across diverse temporal domains and application scenarios. The selected datasets represent critical real-world forecasting challenges spanning energy management, meteorological prediction, and financial markets.
  • Electricity Transformer Temperature (ETT) Datasets: The ETT collection comprises four distinct datasets (ETTh1, ETTh2, ETTm1, and ETTm2) that monitor temperature variations in electricity transformers alongside corresponding load measurements. These datasets facilitate the prediction of future temperature profiles and electrical loads based on historical patterns, which is essential for transformer maintenance and grid stability. The collection offers varied temporal granularities: ETTh1 and ETTh2 provide hourly measurements, while ETTm1 and ETTm2 capture data at 15-minute intervals, enabling comprehensive evaluation across different forecasting horizons and temporal resolutions.
  • Weather: The weather dataset incorporates comprehensive meteorological measurements recorded at 10-minute intervals from a dedicated weather station. This dataset supports forecasting of various atmospheric phenomena, providing crucial information for agricultural planning, transportation safety, and general societal planning activities. The multivariate nature of weather data makes it particularly challenging for forecasting models, as it requires capturing complex interdependencies among atmospheric variables.
  • Electricity (ECL): This dataset encompasses hourly electricity consumption records from 321 individual clients, providing comprehensive insights into consumption patterns and enabling accurate demand forecasting. The dataset is particularly valuable for optimizing power generation scheduling and distribution network management, representing a critical application domain for time series forecasting in energy systems.
  • Exchange Rate: This dataset contains daily exchange rates of the US dollar against eight different currencies, spanning from 1990 to 2016. It represents financial time series forecasting challenges with inherent volatility and complex market dynamics.
Table 4 presents the detailed characteristics of all datasets employed in this study. The collection spans multiple temporal frequencies (10-minute, 15-minute, hourly, and daily intervals) and varies significantly in dimensionality, from 7-dimensional ETT datasets to the 321-dimensional ECL dataset. Dataset sizes are reported in the format (Training, Validation, Test) to clearly delineate the data partitioning strategy employed across all experiments.
A.3 Implementation Details
We summarize the hyperparameter configurations used for EntroPE across all benchmark datasets in Table 6. The table reports dataset-specific settings alongside common defaults, covering patching, embedding, optimization, and training parameters to ensure reproducibility and fair comparison with prior work.
All experiments are conducted with 5 different random seeds, and reported results correspond to the best performance metrics across these runs. Experiments were performed on multiple GPU platforms, including NVIDIA RTX 4090, RTX A5000, NVIDIA GeForce RTX 4090 D, and NVIDIA RTX 6000 Ada-16Q.
Entropy Model Pre-Training. We adopt the exact configuration described in the EDP section 3.3. The model hyperparameters are summarized in Table 5. We pre-train this lightweight model with early stopping, where training is terminated once the validation loss fails to improve by more than 7%. The same setting is applied across all datasets, except for the Electricity dataset where we use a batch size of 32 (batch size of 128 is used for all others).
Table 5: Entropy model configuration used for pre-training.
Hyperparameter Value
Number of layers (n_layern\_layer) 2
Number of heads (n_headn\_head) 4
Embedding dimension (n_embdn\_embd) 8
Dropout 0.1
Bias False
Vocabulary size (vocab_sizevocab\_size) 256
Block size (block_sizeblock\_size) 96
Table 6: Hyperparameter settings for EntroPE across datasets.
Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Rate
dim 8 8 16 16 16 32 8
heads 2 2 2 4 2 4 2
layers 1 2 1 1 2 2 1
max patch length 24 24 24 24 24 24 16
batch_size 64 64 32 32 128 32 32
learning_rate 0.001 0.01 0.01 0.001 0.01 0.01 0.01
dropout 0.05 0.1 0.1 0.1 0.2 0.1 0.2
monotonicity 1 1 1 1 1 1 1
threshold ($\theta$) 3 3.5 3.5 3 3 3.5 3.5
threshold_add ($\gamma$) 0.25 0.2 0.3 0.25 0.35 0.3 0.3
train_epochs 20 20 20 20 20 20 20
patience 7 7 7 7 7 7 7
A.4 Extended Results
A.4.1 Full Results (Long-term forecasting). Input length 96.
We adopted the training criteria established by TimeKAN for fair comparison across all baseline methods. The results for AutoFormer, FEDformer, TimesNet, PatchTST, Time-FFM, iTransformer, and TimeMixer were directly extracted from the TimeKAN paper to ensure consistency in experimental conditions. For HDMixer, TimeBase, CALF, and FilterTS, we executed the original implementations following identical training criteria. Specifically for FilterTS, we addressed the challenge of multiple configuration options by running experiments across all different settings proposed by the authors for each forecasting horizon. We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section A.4.2). LangTime baseline results were obtained directly by running the author’s implementation to maintain consistency with the their reported performance. This methodology ensures that all comparative results are obtained under equivalent experimental conditions, providing a fair and rigorous evaluation framework for our proposed approach. See Tab 7 and Tab. 8 for full results.
Table 7: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model. Look-back window $L$ set to 96 for all cases, forecast window $T=\{96,192,336,720\}$
Dataset T EntroPE TimeKAN HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.327 0.365 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
192 0.360 0.379 0.354 0.380 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
336 0.388 0.399 0.386 0.403 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
720 0.446 0.433 0.452 0.442 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
Avg. 0.378 0.391 0.380 0.398 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
ETTm2 96 0.180 0.265 0.175 0.258 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
192 0.241 0.311 0.241 0.301 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
336 0.313 0.352 0.308 0.349 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
720 0.411 0.413 0.417 0.414 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
Avg. 0.286 0.335 0.285 0.331 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
ETTh1 96 0.375 0.399 0.367 0.395 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
192 0.423 0.425 0.416 0.424 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
336 0.429 0.432 0.446 0.436 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
720 0.439 0.454 0.441 0.453 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
Avg. 0.416 0.425 0.418 0.427 0.448 0.437 0.463 0.429 0.441 0.435 0.440 0.432 0.437 0.425 0.459 0.444 0.442 0.434
ETTh2 96 0.281 0.336 0.290 0.339 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
192 0.371 0.393 0.386 0.399 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
336 0.392 0.394 0.441 0.447 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
720 0.421 0.427 0.450 0.458 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
Avg. 0.366 0.387 0.391 0.410 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
Weather 96 0.164 0.211 0.163 0.209 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
192 0.210 0.252 0.208 0.251 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
336 0.256 0.290 0.263 0.290 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
720 0.339 0.342 0.341 0.341 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
Avg. 0.242 0.273 0.244 0.273 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
Electricity 96 0.163 0.252 0.174 0.266 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
192 0.177 0.268 0.182 0.273 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
336 0.194 0.284 0.197 0.286 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
720 0.235 0.321 0.236 0.320 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
Avg. 0.182 0.271 0.197 0.286 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299
Exchange 96 0.083 0.140 0.084 0.201 0.089 0.201 0.087 0.205 0.081 0.201
192 0.177 0.235 0.173 0.295 0.175 0.298 0.179 0.300 0.168 0.293
336 0.299 0.371 0.316 0.407 0.329 0.409 0.333 0.417 0.299 0.396
720 0.761 0.869 0.827 0.685 0.852 0.690 0.912 0.719 0.805 0.674
Avg. 0.331 0.386 0.355 0.400 0.361 0.400 0.378 0.410 0.338 0.391
Table 8: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model. The look-back window is fixed at $L=96$, with forecast horizons $T=\{96,192,336,720\}$. Baseline results up to 2024 are extracted from Huang et al. (2025b).
Dataset T EntroPE iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
192 0.360 0.379 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
336 0.388 0.399 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
720 0.446 0.433 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
Avg. 0.378 0.391 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
ETTm2 96 0.180 0.265 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
192 0.241 0.311 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
336 0.313 0.352 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
720 0.411 0.413 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
Avg. 0.286 0.335 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
ETTh1 96 0.375 0.399 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
192 0.423 0.425 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
336 0.429 0.432 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
720 0.439 0.454 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
Avg. 0.416 0.425 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
ETTh2 96 0.281 0.336 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
192 0.371 0.393 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
336 0.392 0.394 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
720 0.421 0.427 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
Avg. 0.366 0.387 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
Weather 96 0.164 0.211 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
192 0.210 0.252 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
336 0.256 0.290 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
720 0.339 0.342 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
Avg. 0.242 0.273 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
Electricity 96 0.163 0.252 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
192 0.177 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
336 0.194 0.284 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
720 0.235 0.321 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
Avg. 0.182 0.271 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
Exchange 96 0.083 0.140 0.086 0.206 0.088 0.205 0.107 0.234 0.088 0.218 0.148 0.278 0.197 0.323
192 0.177 0.235 0.177 0.299 0.176 0.299 0.226 0.344 0.176 0.315 0.271 0.315 0.300 0.369
336 0.299 0.371 0.331 0.417 0.301 0.397 0.367 0.448 0.313 0.427 0.460 0.427 0.509 0.524
720 0.761 0.869 0.847 0.691 0.901 0.714 0.964 0.746 0.839 0.695 1.195 0.695 1.447 0.941
Avg. 0.331 0.386 0.360 0.403 0.367 0.404 0.416 0.443 0.354 0.414 0.519 0.429 0.613 0.539
A.4.2 FilterTS Performance Comparison Across Configurations
The original FilterTS results (9) were reported using different hyperparameter settings for different forecast lengths, which complicates direct comparison with other baselines. To ensure fairness, we systematically re-evaluate all provided configurations (Config 1-4) across all forecast horizons and report the configuration that achieves the best performance in terms of MSE and MAE.
Table 9: Comprehensive evaluation of FilterTS across multiple hyperparameter configurations. We assess all available configurations (Config 1-4) for each forecast horizon and report the best-performing results based on MSE and MAE. Missing values (-) denote configurations not originally reported for certain horizons in the official implementation.
Dataset L T Config 1 Config 2 Config 3 Config 4
MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 96 0.321 0.360 0.323 0.362 0.330 0.366
96 192 0.363 0.382 0.364 0.382 0.366 0.383
96 336 0.397 0.405 0.396 0.404 0.398 0.403
96 720 0.478 0.447 0.462 0.441 0.462 0.438
Avg. 0.390 0.399 0.386 0.397 0.389 0.398
ETTm2 96 96 0.174 0.256 0.174 0.257 0.174 0.256
96 192 0.239 0.300 0.238 0.299 0.240 0.299
96 336 0.299 0.338 0.300 0.340 0.301 0.339
96 720 0.405 0.399 0.406 0.399 0.399 0.399
Avg. 0.279 0.323 0.280 0.324 0.279 0.323
ETTh1 96 96 0.375 0.390 0.381 0.396 0.377 0.391
96 192 0.424 0.421 0.431 0.423 0.431 0.423
96 336 0.479 0.451 0.465 0.442 0.479 0.448
96 720 0.480 0.471 0.495 0.481 0.471 0.466
Avg. 0.440 0.433 0.443 0.436 0.440 0.432
ETTh2 96 96 0.289 0.338 0.293 0.344 0.291 0.341
96 192 0.374 0.390 0.374 0.389 0.394 0.400
96 336 0.416 0.426 0.411 0.423 0.433 0.431
96 720 0.421 0.440 0.423 0.441 0.431 0.442
Avg. 0.375 0.399 0.375 0.399 0.387 0.404
Weather 96 96 0.161 0.208 0.180 0.228 0.182 0.228 0.161 0.208
96 192 0.226 0.264 0.210 0.252 0.213 0.255 0.226 0.264
96 336 0.279 0.304 0.286 0.307 0.263 0.292 0.279 0.304
96 720 0.345 0.344 0.349 0.347 0.351 0.348 0.345 0.344
Avg. 0.253 0.280 0.256 0.284 0.252 0.281 0.253 0.280
Electricity 96 96 0.151 0.245 0.153 0.247
96 192 0.164 0.256 0.168 0.260
96 336 0.181 0.274 0.187 0.278
96 720 0.241 0.326 0.227 0.313
Avg. 0.184 0.275 0.184 0.275
Exchange Rate 96 96 0.085 0.201 0.083 0.200 0.084 0.201 0.082 0.198
96 192 0.180 0.300 0.171 0.294 0.173 0.295 0.172 0.294
96 336 0.342 0.420 0.323 0.411 0.316 0.407 0.322 0.410
96 720 0.878 0.699 0.916 0.713 0.827 0.685 0.828 0.686
Avg. 0.371 0.405 0.373 0.405 0.350 0.397 0.351 0.397
A.5 Entropy Qualitative Analysis
Entropy Visualization: Figure 6 demonstrates how entropy-driven boundaries align with temporal transitions, creating patches that preserve intra-patch coherence while capturing natural temporal structure. This validates our core hypothesis that content-aware segmentation outperforms arbitrary fixed-length approaches.
In figure 6,
Panel (a) - Original Time Series: Shows the normalized time series data from the ETTh1 dataset (Sample 0), exhibiting typical temporal patterns with periodic fluctuations and trend changes. The series demonstrates varying levels of predictability across different time segments, ranging from smooth transitions to sharp discontinuities.
Panel (b) - Tokenized Sequence: Displays the discrete token representation of the continuous time series after quantization using the MeanScaleUniformBins tokenizer. Token values range approximately from 50 to 180, capturing the underlying temporal dynamics while enabling discrete sequence modeling. The step-wise pattern reflects the quantization process that maps continuous values to discrete vocabulary tokens.
Panel (c) - Token-wise Entropy with Threshold Regions: Presents the entropy values computed by the pre-trained entropy model for each token position. High entropy regions (red shading) indicate positions where the next token is highly uncertain, suggesting natural breakpoints for dynamic patching. Low entropy regions (green shading) represent predictable sequences that can be grouped into coherent patches. The entropy fluctuates between approximately 3.2-4.4 nats, with clear temporal patterns corresponding to the underlying time series structure.
Panel (d) - Time Series with Entropy-based Coloring: Overlays the original time series with color-coding based on entropy values. Yellow points indicate high entropy (high uncertainty), while dark blue/brown points represent low entropy (high predictability). This visualization reveals the relationship between time series patterns and predictive uncertainty, where rapid transitions and trend changes typically correspond to higher entropy values.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 6: Comprehensive entropy analysis for dynamic patching on ETT and Electricity data. Subfigures show how entropy-driven patches align with temporal transitions.
A.6 Time Series Quantization
Real-valued time series data cannot be directly processed by entropy-based boundary detection methods that rely on discrete probability distributions. Following Ansari et al. (2024), we employ quantization to convert continuous values into discrete tokens while preserving temporal structure.
Given a scaled time series $\tilde{x}_{1:C+H}=[\tilde{x}_{1},\ldots,\tilde{x}_{C},\ldots,\tilde{x}_{C+H}]$, we define a quantization function $q:\mathbb{R}\to\{1,2,\ldots,B\}$ that maps real values to discrete bins. The quantization and dequantization functions are defined as:
$q(x)=\begin{cases}1&\text{if }-\infty\leq x<b_{1},\\ 2&\text{if }b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if }b_{B-1}\leq x<\infty,\end{cases}\quad\text{and}\quad d(j)=c_{j}$ (10)
Unlike traditional approaches that rely on dequantization for prediction, our method uses quantization purely for entropy-based boundary detection while employing linear projection on continuous representations for forecasting.
A.7 Entropy Calculation
We begin with the standard Shannon entropy for a discrete random variable $X$ with possible values $\{x_{1},x_{2},\ldots,x_{n}\}$:
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$ (11)
For sequential data, we are interested in the entropy of a variable $X_{i+1}$ conditioned on the previous observations $X_{\leq i}=\{X_{1},X_{2},\ldots,X_{i}\}$. This leads to the conditional entropy:
$H(X_{i+1}|X_{\leq i})=-\sum_{v\in\mathcal{V}}p(X_{i+1}=v|X_{\leq i})\log p(X_{i+1}=v|X_{\leq i})$ (12)
In practice, we estimate these conditional probabilities using a parameterized model $\theta$, giving us:
$H(x_{i})=-\sum_{v\in\mathcal{V}}p_{\theta}(x_{i+1}=v|x_{\leq i})\log p_{\theta}(x_{i+1}=v|x_{\leq i})$ (13)
where:
  • $H(x_{i})$ represents the conditional entropy at position $i$
  • $p_{\theta}(x_{i+1}=v|x_{\leq i})$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$
  • $x_{\leq i}=\{x_{1},x_{2},\ldots,x_{i}\}$ denotes all observations up to position $i$
  • $\theta$ represents the learnable parameters of the predictive model
This conditional entropy quantifies the uncertainty in predicting the next value given the historical context, with higher values indicating less predictable (more informative) regions of the time series.
A.8 Cross-Entropy Loss
$L_{CE}=-\sum_{i=1}^{N-1}\log P(t_{i+1}|t_{1},\ldots,t_{i};\theta)$ (14)
where $\theta$ represents the model parameters.
A.9 Mean Squared Error (MSE)
$L_{MSE}=\frac{1}{N}\sum_{i=1}^{N}({y_{i}-\hat{y_{i}}})^{2}$
A.10 EntroPE Model Algorithms
Algorithm 1 EntroPE Forward Pass
1:Raw time series $X\in\mathbb{R}^{B\times C\times L}$, forecast horizon $T$
2:Predictions $\hat{Y}\in\mathbb{R}^{B\times C\times T}$
[…]
14:return $\hat{Y}$
Algorithm 2 Entropy-Based Dynamic Patcher (EDP)
1:Tokenized sequence $X^{tok}\in\mathbb{R}^{(B\cdot C)\times L}$, pre-trained entropy model $\theta$
2:Patch boundaries $\mathcal{B}$, patch mask $M\in\mathbb{R}^{(B\cdot C)\times L}$
3:Initialize $\mathcal{B}\leftarrow\emptyset$, $M\leftarrow\mathbf{0}$
4:Phase 1: Entropy Calculation
5:for $t=1$ to L1L-1 do
6:  $\text{context}\leftarrow X^{tok}[:,1:t]$
7:  $\text{logits}\leftarrow\text{entropy\_model}_{\theta}(\text{context})$
8:  $p_{\theta}(x_{t+1}|x_{\leq t})\leftarrow\text{softmax}(\text{logits})$
9:  $H(x_{t})\leftarrow-\sum_{v\in\mathcal{V}}p_{\theta}(v)\log p_{\theta}(v)$
10:end for
11:Phase 2: Boundary Detection
12:for $t=2$ to L1L-1 do
13:  if $H(x_{t})>\theta$ and $H(x_{t})-H(x_{t-1})>\gamma$ then
14:   $\mathcal{B}\leftarrow\mathcal{B}\cup\{t\}$;  $M[:,t]\leftarrow 1$
15:  end if
16:end for
17:Add first and last positions as boundaries
[…]
19:M[:,1]1M[$\leftarrow 1$;  $M[:,L]\leftarrow 1$
20:return $\mathcal{B},M$
Algorithm 3 Adaptive Patch Encoder (APE)
1:Continuous sequence $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$, patch boundaries $\mathcal{B}$, vocab size $V$
2:Patch embeddings $P^{in}$, hidden states $H^{enc}$
[…]
5:Extract variable-length patches
[…]
7:for $i=1$ to $|\mathcal{B}|-1$ do
8:  $\text{patch}_{i}\leftarrow X^{ci}[:,\mathcal{B}[i]:\mathcal{B}[i+1]]$
9:  Append $\text{patch}_{i}$ to patches
10:end for
11:Initial patch embedding via max pooling
12:for each patch $p_{j}$ do
13:  $\text{emb}\leftarrow E(\text{tokenize}(p_{j}))$
14:  $P_{0}[:,j,:]\leftarrow\text{MaxPool}(\text{emb})$
15:end for
16:Cross-attention refinement (NN layers)
[…]
18:for $n=1$ to NN do
19:  $h_{n}\leftarrow\text{TransformerEncoder}_{n}(h_{\text{curr}})$
20:  $Q\leftarrow W_{q}(P_{\text{curr}})$
21:  $K\leftarrow W_{k}(h_{n})$,   $V\leftarrow W_{v}(h_{n})$
22:  $\text{out}\leftarrow\text{softmax}\!\left(\tfrac{QK^{T}}{\sqrt{d_{k}}}\right)V$
23:  $P_{n}\leftarrow P_{\text{curr}}+W_{o}(\text{out})$
24:  Update $h_{\text{curr}},P_{\text{curr}}$
25:end for
26:return $P^{in}\leftarrow P_{N}$,   $H^{enc}\leftarrow h_{N}$
A.11 Use of Large Language Models
Parts of the text in this paper were refined with the assistance of a large language model (ChatGPT, GPT-5), used exclusively for language polishing and improving clarity of exposition. All ideas, methodology, experiments, and analyses were conceived and executed solely by the authors.

### v2

@@
-EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting
+
 Abstract
-Transformer-based models have significantly advanced time series forecasting, with patch-based input strategies offering efficiency and improved long-horizon modeling.
+Patch-based transformers have emerged as efficient and improved long-horizon modeling architectures for time series modeling.
 Yet, existing approaches rely on temporally-agnostic patch construction, where arbitrary starting positions and fixed lengths fracture temporal coherence by splitting natural transitions across boundaries.
 This naive segmentation often disrupts short-term dependencies and weakens representation learning.
-In response, we propose EntroPE (Entropy-Guided Dynamic Patch Encoder), a novel, temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries.
+We propose a novel Entropy-Guided Dynamic Patch Encoder (EntroPE), as a temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries.
 This preserves temporal structure while retaining the computational benefits of patching.
 EntroPE consists of two key modules, namely an Entropy-based Dynamic Patcher (EDP) that applies information-theoretic criteria to locate natural temporal shifts and determine patch boundaries, and an Adaptive Patch Encoder (APE) that employs pooling and cross-attention to capture intra-patch dependencies and produce fixed-size latent representations.
-These embeddings are then processed by a global transformer to model inter-patch dynamics.
-Experiments across long-term forecasting benchmarks demonstrate that EntroPE improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling.
-Code is available at https://github.com/Sachithx/EntroPE.
+Extensive experiments on long-term forecasting, classification, and anomaly detection demonstrate that the proposed method improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling.
 1 Introduction
-Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer & Krauss, 2018), healthcare monitoring (Reis & Mandl, 2003), and environmental surveillance (Smith, 1989).
-However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim & Zohren, 2021; Torres et al., 2021).
-The transformer architecture (Vaswani et al., 2017) has revolutionized sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series forecasting.
+Figure 1: Comparison of patching strategies.
+(a) Temporally-agnostic fixed-length patching places boundaries at arbitrary positions, fragmenting coherent temporal patterns (highlighted regions show split transitions).
+(b) Our temporally-informed approach places boundaries at entropy peaks, where conditional entropy scale indicates natural transition points, preserving temporal coherence within patches.
+Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer and Krauss, 2018), healthcare monitoring (Reis and Mandl, 2003), and environmental surveillance (Smith, 1989).
+However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim and Zohren, 2021; Torres et al., 2021).
+The transformer architecture (Vaswani et al., 2017) has revolutionized language and sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series analysis.
 Despite their promise in learning long-term temporal relationships, time series transformers remain suboptimal due to unique complexities such as non-stationarity, seasonality, varying temporal granularities (Wen et al., 2023; Zeng et al., 2023).
 Recent advances in time series transformers include Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), which address long-sequence modeling challenges through sparse attention mechanisms, decomposition techniques, and improved positional encodings while operating on point-wise inputs.
-Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and long-term forecasting performance.
+Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and performance.
 This demonstrated the critical impact of input tokenization strategies on transformer performance in time series.
 However, subsequent approaches have predominantly adopted similar temporally-agnostic patching schemes, which may not fully capture the temporal coherence and statistical properties inherent in time series data.
-Such temporally-agnostic patching introduces critical technical challenges.
-Fixed-length patching creates inconsistent input representations between training and inference phases.
-During training, time series samples from different time periods result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another).
-However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift where the model encounters patch configurations it rarely or never saw during training.
-Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for underlying dynamics (see Fig.
+These patching strategies treat time as temporally-homogeneous, partitioning sequences at arbitrary positions without regard for underlying temporal structure.
+Temporally-agnostic patching introduces two fundamental challenges.
+They create inconsistent input representations between training and inference phases.
+While training observes patches starting at diverse temporal offsets, result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another).
+However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift in patch configurations.
+Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for frequent underlying dynamics (see Fig.
 1(a)).
-For example, a gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition and forecasting.
-Figure 1: (a) Temporally-agnostic, and (b) Temporally-informed (proposed) patching with calculated entropies for ETTh1 sample.
-These challenges directly impact model performance through two mechanisms.
+For example, a frequently occurring gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition.
+These challenges directly impact model performance.
 Fragmented temporal patterns lead to incomplete intra-patch representations, where semantically related time points are separated across different patches and processed independently.
@@
 This is particularly detrimental for capturing rapid transitions and fine-grained temporal dynamics that require consistent temporal context.
-Unlike prior work that treats patching as a static preprocessing step, we propose EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates dynamic boundary detection as a core architectural component.
-Our approach employs entropy-based criteria to identify natural temporal transition points, ensuring patch boundaries align with the underlying temporal structure rather than arbitrary positions (see Figure 1(b)).
+We argue that patch boundaries should align with points of elevated predictive uncertainty, where the past provides insufficient information to determine the future.
+Formally, such moments are characterized by high conditional entropy of the next observation given historical context.
+Segmenting time at these entropy peaks localizes regime changes at patch boundaries while yielding patches whose interiors remain predictively coherent.
+Motivated by this principle, we introduce EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates uncertainty-aware boundary detection as a core architectural component.
+Unlike prior approaches that fix patch length or rely on heuristic change-point detectors, EntroPE derives segmentation from information-theoretic predictive structure, ensuring patch boundaries align with natural temporal transitions (Fig.
+1(b)).
 We then employ adaptive encoding mechanisms to process these variable-length patches while preserving intra-patch dependencies.
 This temporally-informed approach addresses both the train-inference mismatch and boundary fragmentation issues inherent in temporally-agnostic patching methods.
-The main contributions of this paper are summarized as follows:
-  • We introduce the first information-theoretic approach to patching for time series forecasting, where conditional entropy is used to place boundaries that respect temporal causality and predictive difficulty.
-  • We design an adaptive encoder that converts variable-length patches into fixed representations via pooling and cross-attention, enabling efficient batch computation.
-  • We demonstrate consistent gains across diverse benchmarks, showing that entropy-guided dynamic patching is a practical and generalizable direction for time series transformers.
+The main contributions of this paper are:
+  • We propose EntroPE, the first information-theoretic framework for dynamic patching in time series transformers.
+  • We introduce an adaptive segmentation mechanism that balances modeling accuracy and attention complexity by dynamically controlling patch resolution.
+  • We demonstrate consistent improvements across forecasting, classification, and anomaly detection, establishing entropy-guided dynamic patching as a practical and generalizable paradigm for time series analysis.
+Code is publicly available at https://github.com/Sachithx/EntroPE.
 2 Related Work
-Time Series Transformers and Long-Term Forecasting.
-The transformer architecture has fundamentally transformed sequence modeling across domains, with time series forecasting being no exception.
-Early adaptations like Informer (Zhou et al., 2021) addressed the quadratic complexity of self-attention in long sequences through ProbSparse attention mechanisms, while Autoformer (Wu et al., 2021) and FEDformer (Zhou et al., 2022) introduced decomposition-based approaches and frequency domain modeling to capture complex temporal dependencies.
-These foundational works established transformers as viable architectures for time series tasks but operated on point-wise inputs, leading to computational challenges and a limited ability to capture local temporal patterns efficiently.
-The fundamental challenge these approaches faced was the trade-off between sequence length and computational tractability.
-Point-wise processing, while preserving fine-grained temporal information, resulted in prohibitively long input sequences for practical applications, particularly in long-term forecasting scenarios where historical contexts spanning hundreds or thousands of time steps are crucial for accurate predictions.
-Patch-Based Input Strategies.
-Inspired by vision transformers (Dosovitskiy et al., 2021), patch-based input strategies have been widely adopted in time series forecasting.
-Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) segment time series into fixed-length patches that serve as input tokens.
-This design achieves a quadratic reduction in sequence length while preserving local semantic information through channel-independent processing, demonstrating that tokenization strategies can significantly improve both computational efficiency and forecasting accuracy.
-Building on this foundation, Crossformer (Zhang & Yan, 2023) introduced hierarchical mechanisms tailored for modeling both intra-patch and inter-patch dependencies.
-CARD (Xue et al., 2024) extended patching to multivariate forecasting, while patch-independence approaches (Lee et al., 2024) revealed superior representations in self-supervised settings which further highlights the importance of patch-based input strategies for time series modeling.
-Beyond Transformer architectures, xPatch (Stitsyuk & Choi, 2025) combined dual-flow MLP–CNN designs with exponential decomposition, whereas Pathformer (Chen et al., 2024) and PatchMLP (Tang & Zhang, 2025) leveraged hierarchical processing to capture multi-scale relationships.
-However, fixed-length patching suffers from important limitations.
-Uniform segmentation can fragment coherent patterns such as seasonal transitions or trend shifts.
-During training, windows are sampled by shifting step by step so all possible boundaries are eventually seen, whereas inference fixes a single segmentation that may not align with natural transitions, weakening representation learning and harming generalization.
-In response, MSPatch (Cao et al., 2025) introduces multi-scale patch segmentation to capture both short- and long-term dependencies, mitigating the fragmentation of coherent temporal patterns.
-HDMixer (Huang et al., 2024) proposes a Length-Extendable Patcher (LEP) module that adjusts patch lengths via bi-linear interpolation to enrich boundary information, though it remains constrained by predetermined patch counts and may introduce interpolation-induced uncertainty.
-Beyond patch resizing, foundation models have explored alternative input constructions such as MOIRAI (Woo et al., 2024) leverages frequency-based patch sizing to preserve periodic structures, while Chronos (Ansari et al., 2024) bypasses patching entirely by treating individual time points as tokens, albeit at a substantially higher computational cost.
-While recent works have advanced patch construction strategies, significant gaps remain.
-Existing adaptive methods often overlook causal structure and temporal coherence.
-This prevents them from capturing optimal boundaries guided by temporal uncertainty and predictive difficulty.
-Given the highly dynamic nature of time series across domains, a principled and fully adaptive approach is required.
-Moreover, efficient training capabilities are required, particularly due to the variable number of patches in each sample.
-In contrast, dynamic patching schemes in other domains, such as natural language processing, have shown superior performance and efficiency.
-For instance, the Byte Latent Transformer (Pagnoni et al., 2025) employes byte level inputs instead of traditional word tokenization, and dynamically decide patches as a part of the architeture.
-Motivated by these advances, our work introduces entropy-guided dynamic boundary detection that respects temporal causality and predictive uncertainty, coupled with adaptive encoding mechanisms specifically designed to preserve temporal dependencies within variable-length patches.
-3 Methodology
-3.1 Problem Formulation
-Given a multivariate time series $X=[x_{1},x_{2},\ldots,x_{C}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, the goal is to predict $\hat{Y}\in\mathbb{R}^{C\times T}$ over forecast horizon $T$.
-Following the channel-independence principle from PatchTST (Nie et al., 2023), we deliberately disregard cross-channel dependencies and apply a shared function $f_{\phi}:\mathbb{R}^{L}\to\mathbb{R}^{T}$ (with shared parameters $\phi$) to each channel independently: $\hat{\mathbf{y}}_{c}=f_{\phi}(x_{c}),\quad c=1,2,\ldots,C$.
-This design choice isolates the temporal modeling capabilities from cross-variate relationships, allowing us to focus purely on the effectiveness of our dynamic patching mechanism within each univariate series.
-The multivariate prediction is constructed by stacking channel-wise results:
-$\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{C}]\in\mathbb{R}^{C\times T}$
+Transformers have significantly advanced time series modeling by enabling long-range dependency learning through self-attention.
+Early adaptations such as Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022) improved scalability via sparse attention, decomposition, and frequency-domain modeling.
+However, these approaches operate on point-wise tokens, resulting in prohibitively long sequences and a fundamental trade-off between temporal resolution and computational tractability.
+Patch-Based Input Representations.
+Inspired by Vision Transformers (Dosovitskiy et al., 2021), patch-based methods reduce sequence length while preserving local temporal semantics.
+Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) pioneered fixed-length temporal patching, demonstrating substantial gains in efficiency and forecasting accuracy.
+Subsequent work extended this paradigm through hierarchical dependency modeling (Crossformer (Zhang and Yan, 2023)), multivariate forecasting adaptation (CARD (Xue et al., 2024)), and patch-independent self-supervised learning (Lee et al., 2024).
+Beyond transformers, xPatch (Stitsyuk and Choi, 2025), Pathformer (Chen et al., 2024), and PatchMLP (Tang and Zhang, 2025) explore CNN- and MLP-based hierarchical architectures to capture multi-scale temporal patterns.
+Adaptive and Multi-Granularity Patching.
+Recent efforts relax fixed patch sizes through adaptive or multi-scale segmentation.
+MSPatch (Cao et al., 2025) employs multi-resolution patches, HDMixer (Huang et al., 2024) enables length-extendable interpolation, and MOIRAI (Woo et al., 2024) determines patch size in the frequency domain.
+AdaPatch (Liu et al., 2025a) mitigates non-stationarity via patch-level normalization and reconstruction, but operates over fixed segmentation rather than learning boundary placement.
+APN (Liu et al., 2025c) introduces time-aware adaptive patching for irregular time series through soft window aggregation, emphasizing data-density regularization instead of causality driven segmentation.
+SRSNet (Wu et al., 2025) constructs a selective representation space by scoring, selecting, and reassembling candidate patches, but operates over dense sliding-window patch pools rather than discovering causally meaningful boundary locations.
+TimeMosaic (Ding et al., 2025) further adapts patch granularity based on temporal heterogeneity and motif reuse, optimizing representation efficiency and decoding specialization through segment-wise prompting.
+While these methods improve flexibility, their segmentation criteria remain largely heuristic or representation-driven, prioritizing encoding density, frequency structure, or compression objectives over the underlying predictive or causal structure of the time series.
+Causal and Information-Theoretic Perspectives on Segmentation.
+A complementary line of research argues that segmentation should reflect changes in predictive structure.
+In NLP, Byte Latent Transformer (Pagnoni et al., 2025) demonstrates that architecture-integrated dynamic tokenization improves efficiency by adapting token boundaries to information content.
+However, existing time series methods have not systematically grounded patch boundary selection in information-theoretic measures of predictive uncertainty or temporal causality.
+In contrast to prior work that optimizes patching for representational efficiency or frequency structure, EntroPE places patch boundaries at peaks of conditional predictive entropy.
+This identifies moments where the past becomes insufficient to predict the future, typically corresponding to regime shifts or structural transitions.
+This principled segmentation preserves temporal coherence within patches, localizes uncertainty at boundaries, and eliminates train-inference patch mismatch.
+By coupling entropy-guided dynamic patching with adaptive encoding for variable-length segments, our method introduces a fundamentally new criterion for patch construction based on predictive uncertainty rather than heuristic granularity.
+3 Method
+3.1 Preliminaries: Problem Formulation
+Given a multivariate time series $X=[x^{(1)},\ldots,x^{(C)}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, we address three fundamental tasks, namely forecasting, classification, and anomaly detection.
+Following the channel-independence principle (Nie et al., 2023), our architecture $f_{\phi}$ processes each channel independently to obtain channel-wise representations $\mathbf{z}^{(c)}=f_{\phi}(x^{(c)})$ for $c=1,\ldots,C$.
+Here, $f_{\phi}$ encompasses the complete processing pipeline: entropy-guided dynamic patching, adaptive patch encoding, global transformer, and fusion decoder (detailed in Sec.
+3.3-3.6), producing representations $Z=[\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(C)}]\in\mathbb{R}^{C\times d}$ where $d$ is the final representation dimension.
+$Z$ is projected via linear layer based on specific task.
 Figure 2: Comprehensive architecture of EntroPE.
-The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate forecasting.
-3.2 Forward Process
-The forward process of EntroPE (see Fig.
-2) transforms raw multivariate time series into forecasts through a sequence of interconnected components.
-The raw input multivariate time series $X\in\mathbb{R}^{B\times C\times L}$, where $B$, $C$, and $L$ represent batch size, number of channels (variables), and sequence length respectively, is first processed following the channel independence principle, converting it to $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$ for independent processing of each variable.
-As the next step, to mitigate the statistical difference between training and testing distributions, we integrate the Reversible Instance Normalization layer (Kim et al., 2021) which outputs $X^{Norm.}\in\mathbb{R}^{(B\cdot C)\times L}$.
-Following the Chronos tokenization approach (Ansari et al., 2024) (Appendix A.6), we discretize the continuous input sequence into uniform bins based on a fixed-size vocabulary $V$, producing $X^{tok}$.
-The tokenized sequence $X^{tok}$ is then processed by the EDP (explained in Sec.
-3.3), which generates a patch mask $M\in\mathbb{R}^{B\cdot C\times L}$ indicating optimal patch boundary locations.
-Using this patch mask, the original continuous input $X^{ci}$ is segmented into variable-length patches.
-These patches are fed into the APE (explained in Sec.
-3.4), which generates patch embeddings while aggregating intra-patch dependencies.
-The encoder produces two key outputs, namely, (1) patch embeddings $P^{in}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$, which serve as input embeddings for the global transformer, and (2) encoder hidden states $H^{enc}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ that capture temporal information at the original sequence resolution for subsequent decoding head.
-Here, $\hat{p}$ is not fixed due to the dynamic nature of the patching, $d_{t}$ and $d_{p}$ are time point-wise embedding and patch embedding dimensions, respectively.
-The patch embeddings $P^{in}$ are processed by the Global Transformer, which learns inter-patch long-term dependencies and outputs enriched patch embeddings $P^{out}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$.
-Finally, the Fusion Decoder (FD) combines the global transformer outputs ($P^{out}$) with the encoder hidden states ($H^{enc}$) through cross-attention, flattens the enriched representations, and projects them through a linear head and Instance De-Normalization layer to generate predictions for the future $T$ time steps, producing the final forecast $\hat{Y}\in\mathbb{R}^{B\times C\times T}$.
-3.3 Entropy-Based Dynamic Patcher (EDP)
+The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate predictions.
+3.2 Overall Architecture
+Our approach consists of two distinct phases (Fig.
+2), (1) entropy model pre-training for uncertainty quantification, (2) task-specific training with entropy guided dynamic patching.
+Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
+3.3 Entropy Model Pre-training
 As shown in Fig.
 2(A), our dynamic patch boundary detection leverages a lightweight causal transformer to identify temporal transitions based on predictive uncertainty.
-The process operates in two independent phases, namely model pre-training for temporal pattern learning and entropy-guided patch boundary detection with frozen weights.
-Entropy Model Pre-training.
-Drawing inspiration from the decoder-only transformer architecture and cross-entropy training demonstrated in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens.
-Our lightweight architecture comprises 8 embedding dimensions (shared input and head), 2 transformer layers, 4 attention heads, totaling 3,648 learnable parameters.
-The model performs autoregressive prediction at time $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence at time $\{t_{1},t_{2},\ldots,t_{L-1}\}$ using standard cross-entropy loss.
-This streamlined design efficiently captures temporal dynamics while maintaining computational tractability for entropy estimation.
-Crucially, our objective is not to achieve superior forecasting performance, but rather to align the causal transformer for measuring predictive uncertainty across the discrete vocabulary.
-This architectural approach enables identification of natural transition points in raw time series sequences, where predictive uncertainty that quantified through entropy over the output next-token distribution reaches local maxima.
-Entropy-Guided Patch Boundary Detection.
-Upon training convergence, we freeze model weights and compute entropy for each time position.
-For position $t$, we calculate the Shannon entropy of the next-token prediction distribution:
-$H(x_{t})=-\sum_{v\in\mathcal{V}}p_{\theta}(x_{t+1}=v|x_{\leq t})\log p_{\theta}(x_{t+1}=v|x_{\leq t})$ (1)
-Following Pagnoni et al.
-(2025), we implement dual-threshold boundary detection combining global and relative uncertainty measures.
-A patch boundary is placed at position $t$ when three conditions are met simultaneously:
-$H(x_{t})$ $>\theta\quad\text{(global entropy threshold)}$ (2)
-$H(x_{t})-H(x_{t-1})$ $>\gamma\quad\text{(relative entropy increase)}$ (3)
-This mechanism ensures patches begin only at positions with both high absolute uncertainty and significant uncertainty increases, avoiding fragmentation from minor entropy fluctuations while capturing meaningful temporal transitions.
-EDP outputs patch mask $M$, which indicates the non-overlapping patch segmentation for further processing.
+Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens.
+For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq.
+38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
+A critical distinction.
+Unlike Chronos (Ansari et al., 2024), which operates entirely in the quantized domain and suffers from range limitations during dequantization, we use quantization only for computing entropies and input projection.
+The downstream prediction heads project the output-embeddings, not quantized tokens.
+This decoupling eliminates the quantization range constraint as our predictions are not limited to the discrete vocabulary.
+Quantization serves purely as a preprocessing step for the entropy model (Sec.
+3.3), while the actual time series transformer and prediction heads operate on continuous representations throughout.
+Determining Quantization Range.
+To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics.
+Let $X_{\text{train}}$ denote z-score normalized training values.
+We compute empirical quantiles:
+$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (1)
+where $\epsilon=0.005$ captures 99.5% of the distribution.
+The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.
+Model Selection.
+Drawing inspiration from decoder-only transformers in time series (Das et al., 2024) and cross-entropy guided training in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens.
+Our lightweight architecture comprises 8-16 embedding dimensions and 2 transformer layers, totaling $\approx 10$k learnable parameters.
+The model performs autoregressive prediction at timesteps $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence $\{t_{1},t_{2},\ldots,t_{L-1}\}$.
+Model Training.
+We train our lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction.
+The model is optimized via:
+$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (2)
+where $\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})$ is the cross-entropy loss (Eq. 21).
+Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation.
+This decoupling is critical, as the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
+3.4 Entropy-Guided Dynamic Patcher (EDP)
+Entropy Computation.
+For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently.
+For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
+$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=\\ -\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (3)
+yielding entropy sequence,
+$\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}.$ (4)
+As established in Appendix A.2, high conditional entropy indicates positions where the future is less predictable from the past, typically at regime shifts, trend reversals, or volatility bursts.
+By placing boundaries at such points (Proposition A.2), patches become internally coherent while transition uncertainty is localized at edges.
+Adaptive Threshold Determination.
+Rather than using fixed global thresholds requiring manual tuning, we employ sample-adaptive thresholds derived from quantile-based estimation.
+For each input sequence with entropies $\{H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})\}$, we compute sample-specific thresholds.
+Hence, the threshold selection becomes interpretable.
+The absolute entropy threshold is the $\alpha$-th quantile of the sequence’s entropy distribution:
+$\theta_{\text{abs}}^{(c)}=Q_{\alpha}(\mathbf{H}^{(c)}).$ (5)
+The relative threshold is the $\alpha$-th quantile of consecutive differences $\Delta\mathbf{H}^{(c)}=\bigl[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})\bigr]_{t=2}^{L}$,
+$\gamma_{\mathrm{rel}}^{(c)}$ $=Q_{\alpha}\!\left(\Delta\mathbf{H}^{(c)}\right),$ (6)
+where $\alpha$ is the quantile parameter (e.g., $\alpha=75\%$). This formulation enables automatic adaptation to varying entropy levels across sequences, eliminating dataset-specific threshold tuning.
+Boundary Detection.
+A patch boundary is placed at position $t$ based on the following conditions:
+$H(x_{t}^{(c)})$ $>\theta_{\text{abs}}^{(c)}\quad\text{(high absolute uncertainty)}$ (7)
+$\Delta H(x_{t}^{(c)})$ $>\gamma_{\text{rel}}^{(c)}\quad\text{(significant uncertainty increase)}$ (8)
+$M_{t-1}^{(c)}$ $=0\quad\text{(no consecutive boundaries)}.$ (9)
+This threshold mechanism produces a binary patch boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch start.
+The framework supports using either threshold independently or in combination, allowing flexible adaptation to different temporal characteristics.
+EDP outputs mask $M$ for subsequent non-overlapping patch processing.
+Patch Segmentation.
+Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$.
+The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
+$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},$ (10)
+where $\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$, $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$.
+Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
 3.5 Adaptive Patch Encoder (APE)
-The EDP produces variable-length patches where both patch count $\hat{p}$ and individual lengths vary across samples (patch mask $M$).
+The EDP produces variable-length patches where both patch count and individual lengths vary across samples.
 As shown in Fig.
@@
 First initial dimensionality reduction via pooling (from time-point embeddings to initial patch embeddings), followed by iterative cross-attention refinement.
-This design is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), but specialized for temporal patch aggregation.
-Initial Patch Embedding.
-We first obtain the time-point embedding $h_{0}$ for the entire sequence with length $L$ through an Embedding Layer $E\in\mathbb{R}^{V\times d_{t}}$ over the same Vocabulary and Tokenizer used in the Entropy Model.
-For each variable-length patch $p_{j}$, we generate initial patch representations via max pooling: $P_{0,j}=\text{MaxPool}(E_{f}(p_{j}))$, where $E_{f}$ represents the time-point embedding function.
-Max pooling captures salient temporal features while providing translation invariance within patches.
-Cross-Attention Refinement. We refine patch representations using cross-attention, where patch embeddings serve as queries and time-point embeddings act as keys and values. For layer $n$:
-$P_{n}=P_{n-1}+W_{o}\left(\text{softmax}\!\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right),$ (4)
-with attention components:
-$Q_{j}$ $=W_{q}(P_{n-1,j})\quad\text{(patch embeddings as queries)},$ (5)
-$K_{i}$ $=W_{k}(h_{n-1,i}),\quad V_{i}=W_{v}(h_{n-1,i})\quad\text{(time-point embeddings as keys/values)},$ (6)
-where $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ are learnable projection matrices corresponding to queries, keys, values, and output transformations, respectively.
-For $N$-layer architectures with $N>1$, time-point embeddings are iteratively refined through stacked transformer layers as $h_{n}=\text{APETransformer}(h_{n-1})$, with cross-attention mechanisms operating between consecutive layers.
-Here, $h_{n}$ is the time point embedding sequence at layer $n$.
-To ensure that self-attention within each APETransformer operates exclusively within patch boundaries, we employ a tailored attention mask that accommodates the variable patch lengths throughout the sequence.
-This masking strategy prevents information leakage across patch boundaries while preserving intra-patch temporal dependencies.
-The encoder outputs fixed-size patch embeddings $P^{in}=P_{N}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$ for global transformer processing and preserves encoder hidden states $H^{enc}=h_{N}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ for temporal information in the decoding stage.
+This design choice is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), due to their demonstrated performance on language and image data.
+Patch Encoding.
+Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the APE.
+First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
+$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L.$ (11)
+For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
+$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}.$ (12)
+We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
+$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (13)
+where attention weights are computed only within the patch:
+$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (14)
+This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
+Batch Processing with Variable Patches.
+For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$.
+We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
 3.6 Global Transformer & Fusion Decoder
-The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate forecasts.
+The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate predictions.
 Global Transformer.
-The patch embeddings $P^{in}$ from the adaptive encoder are processed through a standard transformer architecture to capture long-range temporal relationships between patches.
-Since intra-patch dependencies are already modeled by the encoder, the global transformer focuses exclusively on inter-patch interactions.
-We employ bidirectional attention (non-causal) to enable comprehensive context modeling across the entire sequence: $P^{out}=\text{GlobalTransformer}(P^{in})$.
-Fusion Decoder (FD).
-A key challenge in dynamic patching is that the number of patches $\hat{p}$ varies across samples, making direct forecasting from flattened patch representations problematic.
-To address this, we employ a cross-attention mechanism that fuses global patch context with fine-grained temporal information preserved from the encoder (see Fig.
-2(C)).
-The decoder reverses the attention configuration used in the adaptive encoder.
-Here, encoder hidden states $H^{enc}$ serve as queries, while global patch embeddings act as keys and values:
-$Q_{i}$ $=W_{q}(H^{enc}_{i})\quad\text{(time-point queries)},$ (7)
-$K_{j}$ $=W_{k}(P^{out}_{j}),\quad V_{j}=W_{v}(P^{out}_{j})\quad\text{(patch keys/values)}$ (8)
-The cross-attention operation enriches time-point representations with global context:
-$H^{dec}=H^{enc}+W_{o}\left(\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\right).$ (9)
-This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, ensuring preservation of both local temporal patterns and global contextual dependencies.
-Optionally, similar to APE, the iterative cross-attention refinement can be configured through alternating transformer layers and cross-attention layers, creating a structured information exchange between different representational levels.
-[…]
+Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
+$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (15)
+Fusion Decoder.
+We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
+$\tilde{H}^{(c)}=H^{(c)}+\\ W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\\ \in\mathbb{R}^{L\times d_{t}}$ (16)
+This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
+Task-Specific Head.
+The enriched representations $\tilde{H}^{(c)}$ are projected to generate future predictions.
+We apply flattening followed by a linear projection with Instance De-Normalization: $\text{InstanceDeNorm}\!\left(W_{\text{proj}}(\text{Flatten}(\tilde{H}^{(c)}))\right)$.
+Training criteria and additional details on task-specific heads are provided in Appendix B.
 4 Experiments
+In the section we evaluate the efficacy of EntroPE on long-term forecasting, classification and anomaly detection tasks.
+We show that our model can serve as a foundation model with competitive performance accross these tasks.
+4.1 Long-Term Forecasting
 Datasets.
-We evaluate EntroPE on seven widely-used long-term multivariate forecasting benchmarks: ETT family (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, and Exchange Rate datasets, covering diverse domains and temporal characteristics (detailed descriptions in Appendix Tab.
-4).
-Baselines.
-We benchmark against 14 state-of-the-art deep forecasting models from 2021–2025, organized by architecture.
-(1) Transformer-based models: iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Kernel-attention model: TimeKAN (Huang et al., 2025b); (5) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (6) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025); (7) Filter-based model: FilterTS (Wang et al., 2025).
-(Further details on baselines are discussed in Appendix A.1.)
-Experimental Settings.
-Following standard practices in time series forecasting literature, we evaluate performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics across multiple prediction horizons $T=\{96,192,336,720\}$ and fixed look-back window $L=96$.
-Adopting the same experiemnt setup as Huang et al.
-(2025b), baseline results up to 2024 are taken from TimeKAN, while other methods are re-evaluated using their official implementations.
-EntroPE experiemnts are conductd in 5 random seeds and we report the best results.
-(see Appendix A.3.)
-4.1 Main Results
-Table 1 summarizes average performance across all prediction horizons.
-EntroPE matches or surpasses, in many cases, recent variable-mixing, Transformer-based, CNN-based, MLP-based, foundation, lightweight, LLM-aligned, filter-based, and KAN-based models under the standard 96-step look-back with horizons up to 720.
-It retains PatchTST’s overall architecture but introduces two key innovations namely, (i) entropy-driven dynamic patching and (ii) adaptive embeddings.
-Relative to PatchTST, EntroPE achieves notable accuracy gains of approximately 20% on ETTh1, 15% on Electricity, and about 10% on average, while also reducing token count and improving efficiency by leveraging channel independence and non-overlapping patches.
-Beyond PatchTST, EntroPE consistently outperforms HDMixer’s length-extendable patching and improves upon TimeBase, though the latter remains highly efficient.
-On the Electricity dataset, iTransformer continues to offer the strongest performance-efficiency tradeoff given its channel-wise self-attention across 321 and 862 variables, respectively.
-CALF and LangTime also remain competitive, though their large model sizes (CALF has 18 million, LangTime has 500 million while EntroPE has 0.1 million range parameters) impose substantial costs.
-Across other benchmarks, EntroPE delivers robust improvements, underscoring the value of entropy-guided temporal segmentation for multivariate forecasting.
+To examine forecasting performance we conduct comprehensive evaluations on six widely-used long-term multivariate forecasting datasets.
+i.e., Electricity (ECL) featuring electricity consumption data, ETT family (ETTh1, ETTh2, ETTm1, ETTm2) that encompass a range of scenarios in energy transfer technology, Weather that offers insights into various meteorological variables over time.
+Detailed descriptions in Appendix Tab.
+6.
+Baselines and Experimental Settings.
+We benchmark against fifteen state-of-the-art baselines.
+(1) Transformer-based models: TimeMosaic (Ding et al., 2025), iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023), TSLANet (Eldele et al., 2024); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (5) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025b); (6) Filter-based model: FilterTS (Wang et al., 2025).
+(Further details on baselines are discussed in Appendix D.1.)
 Table 1: Multivariate time series forecasting results on benchmark datasets.
 Results are averaged across prediction horizons $T=\{96,192,336,720\}$ with fixed input length $L=96$ for all datasets.
-Best results are highlighted in red and second-best in blue.
-A dash (-) indicates that the configuration was not found in the original implementation.
-(Full table: Appendix A.4)
-Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338 0.613 0.539
-FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327 0.519 0.429
-DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319 0.354 0.414
-TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304 0.416 0.443
-PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318 0.367 0.404
-Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299 0.338 0.391
-HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295 - -
-iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270 0.360 0.403
-TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272 0.378 0.410
-TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296 - -
-LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285 0.361 0.400
-TimeKAN [2025] 0.418 0.427 0.391 0.410 0.380 0.398 0.285 0.331 0.244 0.273 0.197 0.286 - -
-FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275 0.355 0.400
-CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266 - -
-EntroPE [2025] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271 0.331 0.386
-4.2 Ablation Studies
+Best results are highlighted in red and second/-best in blue.
+Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
+MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338
+FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327
+DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319
+TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304
+PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
+Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299
+HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295
+iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270
+TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272
+TSLANet [2024] 0.448 0.441 0.372 0.399 0.378 0.397 0.283 0.327 0.259 0.280 0.199 0.283
+TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296
+LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285
+FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275
+CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266
+TimeMosaic [2026] 0.427 0.442 0.369 0.399 0.388 0.392 0.278 0.319 0.253 0.277 0.188 0.281
+EntroPE [2026] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
+Results.
+Table 1 shows EntroPE achieves competitive or superior performance across all baseline categories (Transformer, CNN, MLP, foundation, LLM-aligned, and filter-based models).
+While retaining PatchTST-like architecture, EntroPE introduces dynamic patching and adaptive encoding, yielding  20% MSE reduction on ETTh1,  15% on Electricity, and  10% average improvement over PatchTST, alongside reduced token count through non-overlapping patches.
+Notably, iTransformer and CALF demonstrates remarkable performance on high-dimensional Electricity dataset (321 variables) due to its channel-wise attention design.
+While CALF ($\sim$18M parameters) and LangTime ($\sim$500M parameters) show competitive results, EntroPE achieves comparable accuracy with 500-1000× fewer parameters (100k-1M), demonstrating that entropy-guided segmentation provides robust improvements without requiring massive model capacity.
+Compared to some noteworthy dynamic or adaptive patching approaches such as HDMixer and TimeMosaic, EntroPE consistently delivers stronger or more stable gains across datasets.
+Notably, TimeMosaic employs substantially larger models (3–13M parameters) and higher training durations, whereas EntroPE achieves comparable or better accuracy with smaller models.
+This highlights that performance gains arise from principled entropy-driven boundary placement rather than increased model capacity.
+(Full table: Tab.
+16 & 17)
+4.2 Classification
+Datasets.
+To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA multivariate time series classification datasets (Bagnall et al., 2018).
+We replace the forecasting head with a classification head while keeping the core architecture unchanged.
+Baselines and Experimental Settings.
+We select eight state-of-the-art baselines, i.e., TSLANet, GPT4TS (Zhou et al., 2023), TimesNet, ROCKET (Dempster et al., 2020), TS-TCC (Eldele et al., 2021), Crossformer (Zhang and Yan, 2023) and PatchTST as they showed the best classification accuracy over other Transformer-based architectures.
+Last, we experiment with a simple single-layer MLP.
+Table 2: Classification accuracy (%) on 10 UEA multivariate classification datasets.
+Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
+[2026] [2024] [2023] [2023] [2020] [2023] [2023] [-] [2021]
+EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
+FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
+Handwriting 57.4 57.88 34.56 26.18 48.47 26.24 26.00 22.47 47.76
+Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
+JapaneseVowels 99.2 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
+PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
+SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
+SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
+SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
+UWaveGestureLibrary 94.4 91.25 84.38 83.13 76.03 81.56 80.00 81.88 86.25
+Average 77.13 76.03 67.45 69.36 72.90 70.88 71.14 69.51 73.53
+Results.
+Table 2 shows EntroPE achieves an average accuracy of 77.13%, competitive with TSLANet (76.03%), and outperforming TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines.
+This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
+Anomaly Detection Evaluation.
+Appendix C shows the detailed information on datasets, baselines and results of EntroPE’s Anomaly detection performance.
+4.3 Ablation Study
 Model component analysis.
-To systematically evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720 time steps, measuring performance via MSE.
+To evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720, measuring performance via MSE.
 We progressively remove components from our full EntroPE architecture to create four configurations: (1) EntroPE (Full): Dynamic patching + Adaptive encoder + Fusion decoder; (2) EntroPE - Dynamic: Static patching + Adaptive encoder + Fusion decoder; (3) EntroPE - (Dynamic Patching + Adaptive Encoder): Static patching + Max pooling + Fusion decoder; (4) EntroPE - (Dynamic Patching + Adaptive Encoder + Fusion Decoder): Static patching + Max pooling + Flattened output.
+Table 3: Component importance and dynamic vs. static patching comparison. Best and second-best are highlighted in red and blue.
+Dataset T Full -EDP -EDP -APE -EDP -APE -FD
+Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat
+ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519
+720 0.439 0.460 0.469 0.477 0.461 0.527
+ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462
+720 0.397 0.448 0.470 0.463 0.466 0.456
+ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427
+720 0.445 0.451 0.459 0.452 0.457 0.483
+Weather 336 0.258 0.261 0.265 0.270 0.262 0.262
+720 0.341 0.343 0.353 0.359 0.347 0.346
 The full EntroPE architecture achieves the best performance across all datasets and horizons, while systematic component removal leads to progressive performance degradation (Table 3).
@@
 This systematic performance decline confirms that each architectural choice addresses specific modeling challenges in time series forecasting.
-Table 2: Ablation study showing component importance and dynamic vs. static patching comparison. Best results are highlighted in red and second-best in blue.
-Dataset T Full -EDP -EDP -APE -EDP -APE -FD Baselines
-Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat PatchTST iTrans.
-ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519 0.501 0.487
-720 0.439 0.460 0.469 0.477 0.461 0.527 0.503 0.500
-ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462 0.427 0.428
-720 0.397 0.448 0.470 0.463 0.466 0.456 0.436 0.427
-ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427 0.421 0.426
-720 0.445 0.451 0.459 0.452 0.457 0.483 0.462 0.491
-Weather 336 0.258 0.261 0.265 0.270 0.262 0.262 0.284 0.278
-720 0.341 0.343 0.353 0.359 0.347 0.346 0.356 0.358
 Dynamic vs.
@@
 Experiments are conducted on ETTh1, ETTm1, and Weather datasets with a fixed input length of 96 and forecasting horizons of 336 and 720.
-Dynamic patching achieves the lowest MSE in most settings (Table 2, while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme.
+Dynamic patching achieves the lowest MSE in most settings (Table 3), while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme.
 However, the computational cost of this setting is substantially higher, as the self-attention mechanism must attend to every time step individually.
 Figure 3 illustrates the efficiency advantage of dynamic patching over static alternatives under different configurations.
-Figure 3: Ablation study for Dynamic and Static(patch_length) patching with single epoch duration.
+Figure 3: Dynamic and Static(patch_length) patching with epoch duration.
 […]
-Figure 4: Ablation study for threshold sensitivity with epoch duration.
+Figure 4: Threshold sensitivity with epoch duration.
 4.4 Efficiency Analysis
-Figure 5: EntroPE efficiency analysis.
 As shown in Figures 3 and 4, our model adapts to different patching schemes, directly affecting computational efficiency.
-In the dynamic setting, the model maintains superior predictive performance while operating within an optimal computation budget.
-We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), Mean Squared Error (MSE), and the number of trainable parameters.
-Figure 5 demonstrates that our model achieves a more favorable performance–efficiency trade-off, with bubble size proportional to the parameter count.
-Notably, while TimeMixer and TimeKAN occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture.
+We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), MSE, and the number of trainable parameters.
+Figure 5 demonstrates that our model achieves a more favorable performance-efficiency trade-off, with bubble size proportional to the parameter count.
+Notably, while TimeMixer and TimeKAN (Huang et al., 2025b) occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture.
+(See full table: Tab.
+13)
+Figure 5: EntroPE’s efficiency analysis on ETTm1 dataset.
 5 Conclusion
-We introduced EntroPE, a temporally informed dynamic patch encoding framework that extends time series transformer architectures through entropy-guided boundary detection.
-Instead of partitioning sequences at arbitrary intervals, EntroPE strategically places patch boundaries at natural temporal transitions, preserving temporal coherence while supporting efficient variable-length patch processing.
-This design provides a principled way to incorporate information-theoretic signals into patching, yielding improved forecasting accuracy alongside practical computational benefits.
-Beyond its empirical performance, EntroPE highlights the importance of respecting the intrinsic structure of time series, opening new research directions in adaptive sequence modeling and context-aware temporal representation learning.
-Appendix A Appendix
-A.1 Baseline Details
+We introduced EntroPE, a dynamic patch encoding framework that uses entropy-guided boundary detection to identify natural temporal transitions in time series.
+By placing patch boundaries at information-rich transitions rather than fixed intervals, EntroPE preserves temporal coherence while enabling efficient variable-length processing.
+This information-theoretic approach improves performance across downstream tasks with practical computational benefits.
+Beyond empirical gains, EntroPE demonstrates the value of respecting intrinsic temporal structure, opening new directions in adaptive context-aware sequence modeling.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Preliminaries & Motivation
+A.1 Method Overview
+Our approach consists of two distinct phases: (1) entropy model pre-training for uncertainty quantification, and (2) task-specific training with entropy-guided dynamic patching.
+Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
+A.2 Theoretical Motivation
+We provide a theoretical justification for entropy-guided patch segmentation.
+Our framework rests on two principles: (i) conditional entropy characterizes irreducible predictive uncertainty, and (ii) segmenting sequences at high-entropy points yields patches with lower intra-segment uncertainty and improved temporal coherence.
+Proposition A.1 (Conditional Entropy Lower-Bounds Irreducible Prediction Loss).
+Let $\tau_{1:L}$ denote quantized time-series tokens and let $p_{\star}(\tau_{t+1}\mid\tau_{\leq t})$ be the true next-token conditional distribution.
+For any predictive model $q(\tau_{t+1}\mid\tau_{\leq t})$, the expected one-step negative log-likelihood satisfies
+$\mathbb{E}_{p_{\star}}\!\left[-\log q(\tau_{t+1}\mid\tau_{\leq t})\right]\;\geq\;H_{\star}(\tau_{t+1}\mid\tau_{\leq t}),$ (17)
+with equality if and only if $q(\cdot\mid\tau_{\leq t})=p_{\star}(\cdot\mid\tau_{\leq t})$ almost surely.
+Proof sketch.
+By the cross-entropy decomposition,
+$\mathbb{E}_{p_{\star}}[-\log q]=H_{\star}(\tau_{t+1}\mid\tau_{\leq t})+\mathrm{KL}\!\left(p_{\star}(\cdot\mid\tau_{\leq t})\;\|\;q(\cdot\mid\tau_{\leq t})\right),$
+and since $\mathrm{KL}(\cdot\|\cdot)\geq 0$, the bound follows, with equality if and only if $q=p_{\star}$ almost surely. ∎
+Proposition A.2 (High-Entropy Boundaries Reduce Intra-Patch Predictive Uncertainty).
+Let a segmentation $\mathcal{S}=\{1=b_{1}<b_{2}<\cdots<b_{P+1}=L+1\}$ partition a sequence into patches $b_{j},b_{j+1})$. Define the total within-patch uncertainty mass as
+$U(\mathcal{S})=\sum_{j=1}^{P}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t}).$ (18)
+Consider a candidate boundary $b$ within a patch $[s,e)$, and suppose $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is a local maximum or exhibits a sharp positive increment relative to its neighborhood.
+Then introducing a boundary at $b$ yields a segmentation $\mathcal{S}^{\prime}$ such that the average within-patch entropy
+$\frac{1}{|[b_{j},b_{j+1})|}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$
+is no greater in at least one resulting patch and strictly lower whenever $H_{\star}$ exhibits a concentrated spike at $b$.
+Proof sketch.
+Let $[s,e)$ contain $b\in(s,e)$.
+Splitting into $[s,b)$ and $[b,e)$ preserves total uncertainty mass, but redistributes it across two segments.
+If $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is locally elevated, then keeping $b$ interior inflates the mean uncertainty of the full segment.
+By relocating $b$ to a boundary, the high-entropy region is isolated to the segment edge, lowering the average entropy within at least one subsegment.
+This reduces expected intra-patch predictive uncertainty while preserving segmentation budget.
+∎
+Connection to EntroPE.
+Proposition A.1 establishes conditional entropy as a fundamental lower bound on achievable predictive loss, implying that peaks in entropy correspond to irreducible uncertainty in temporal evolution.
+Proposition A.2 further motivates entropy-guided segmentation by showing that placing boundaries at high-entropy or sharply rising entropy points yields patches whose interiors exhibit lower average predictive uncertainty and greater temporal coherence.
+EntroPE operationalizes this principle by training a lightweight causal transformer on quantized time series, freezing it, and computing per-timestep conditional entropy from the predicted token distribution.
+Since the true entropy $H_{\star}$ is unknown, we approximate it using the frozen model’s predictive distribution, $H_{t}\approx H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$, and apply quantile-based thresholding (Eq.
+56) to construct a boundary mask for dynamic patching.
+Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
+Remark A.3 (Quantization Does Not Limit Predictions).
+While we use quantization for entropy computation, the downstream prediction heads operate on continuous representations (Sec.
+3.6).
+This decoupling eliminates range constraints: predictions are not limited to the discrete vocabulary $\mathcal{V}$ used for boundary detection.
+This theoretical framework justifies our design: EntroPE identifies natural transition points where predictive uncertainty is high, ensuring patches are internally coherent while boundaries mark regime changes.
+A.2.1 Phase I: Entropy Model Pre-training
+Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens.
+For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq.
+38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
+Automatic Quantization Range.
+To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics.
+Let $X_{\text{train}}$ denote z-score normalized training values.
+We compute empirical quantiles:
+$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (19)
+where $\epsilon=0.005$ captures 99.5% of the distribution.
+The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.
+Model Training.
+We train a lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction.
+The model is optimized via:
+$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (20)
+where the cross-entropy loss for a sequence is:
+$\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})=-\frac{1}{L-1}\sum_{t=1}^{L-1}\log p_{\theta_{\text{ent}}}(\tau_{t+1}|\tau_{\leq t})$ (21)
+Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation.
+This decoupling is critical: the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
+A.2.2 Phase II: Entropy-Guided Dynamic Segmentation
+Entropy Computation.
+For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently.
+For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
+$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=-\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (22)
+yielding entropy sequence $\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}$.
+Theoretical Motivation.
+High conditional entropy $H(x_{t}|x_{<t})$ indicates temporal transitions where $x_{t}$ is less predictable from history.
+By placing boundaries at high-uncertainty regions, patches become internally coherent (low intra-patch entropy) while boundaries mark regime changes.
+Unlike fixed-length patching, our proposed method doesn’t ignore predictive structure.
+Adaptive Threshold Determination.
+Rather than fixed thresholds requiring dataset-specific tuning in general, we introduce sample-adaptive thresholds via quantile estimation.
+For the $\alpha$-th quantile (e.g., $\alpha=0.75$), we derive:
+$\theta_{\text{abs}}^{(c)}$ $=Q_{\alpha}(\mathbf{H}^{(c)})\quad\text{(absolute entropy threshold)}$ (23)
+$\Delta\mathbf{H}^{(c)}$ $=[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})]_{t=2}^{L}\quad\text{(entropy differences)}$ (24)
+$\gamma_{\text{rel}}^{(c)}$ $=Q_{\alpha}(\Delta\mathbf{H}^{(c)})\quad\text{(relative entropy threshold)}$ (25)
+Boundary Detection.
+We generate a binary boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch boundary at position $t$. Position $t$ becomes a boundary when:
+$M_{t}^{(c)}=\mathbb{1}\left[H(x_{t}^{(c)})>\theta_{\text{abs}}^{(c)}\land\Delta H(x_{t}^{(c)})>\gamma_{\text{rel}}^{(c)}\land M_{t-1}^{(c)}=0\right]$ (26)
+with $M_{1}^{(c)}=1$ (first position always starts a patch).
+Patch Segmentation.
+Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$.
+The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
+$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$ (27)
+where $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$.
+A.2.3 Phase III: Task-Specific Training with Fixed Boundaries
+Patch Encoding.
+Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the Adaptive Patch Encoder (APE).
+First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
+$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L$ (28)
+For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
+$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}$ (29)
+We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
+$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (30)
+where attention weights are computed only within the patch:
+$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (31)
+This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
+Batch Processing with Variable Patches.
+For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$.
+We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
+Global Transformer.
+Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
+$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (32)
+Fusion Decoder.
+We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
+$\tilde{H}^{(c)}=H^{(c)}+W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\in\mathbb{R}^{L\times d_{t}}$ (33)
+This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
+Appendix B Task-Specific Linear Projection
+Task-Specific Heads.
+The enriched representations are projected to task-specific outputs:
+Forecasting: Channel-wise linear projection
+$\hat{Y}^{(c)}=g_{\text{fc}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{T},\quad c=1,\ldots,C$ (34)
+with MSE loss: $\mathcal{L}_{\text{forecast}}=\frac{1}{C\cdot T}\sum_{c=1}^{C}\sum_{t=1}^{T}(Y_{c,t}^{(c)}-\hat{Y}_{c,t}^{(c)})^{2}$
+Classification: Flatten all channels and project to class logits
+$\hat{y}=g_{\text{cls}}\left(\text{Flatten}([\tilde{H}^{(1)},\ldots,\tilde{H}^{(C)}])\right)\in\mathbb{R}^{K}$ (35)
+with cross-entropy loss: $\mathcal{L}_{\text{class}}=-\log\frac{\exp(\hat{y}_{y})}{\sum_{k=1}^{K}\exp(\hat{y}_{k})}$ where $y\in\{1,\ldots,K\}$ is the true label.
+Anomaly Detection: Channel-wise reconstruction
+$\hat{X}^{(c)}=g_{\text{rec}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{L},\quad c=1,\ldots,C$ (36)
+with MSE loss: $\mathcal{L}_{\text{anomaly}}=\frac{1}{C\cdot L}\sum_{c=1}^{C}\sum_{t=1}^{L}(X_{c,t}^{(c)}-\hat{X}_{c,t}^{(c)})^{2}$
+Training Protocol.
+During task-specific training, the entropy model $\theta_{\text{ent}}^{*}$ remains frozen and is used only to compute patch boundaries via Eq.
+(8)-(10).
+Only the encoder $E$, APE, transformer, fusion decoder, and task head parameters $\Phi=\{E,\theta_{\text{APE}},\theta_{\text{trans}},\theta_{\text{FD}},\theta_{\text{task}}\}$ are optimized:
+$\Phi^{*}=\arg\min_{\Phi}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{\text{task}}}\left[\mathcal{L}_{\text{task}}(X,Y;\Phi\mid\theta_{\text{ent}}^{*})\right]$ (37)
+where $\mathcal{D}_{\text{task}}$ is the task-specific dataset and $\mathcal{L}_{\text{task}}\in\{\mathcal{L}_{\text{forecast}},\mathcal{L}_{\text{class}},\mathcal{L}_{\text{anomaly}}\}$.
+Computational Complexity.
+Let $L$ be sequence length, $P$ the average number of patches, and $d$ the hidden dimension.
+EntroPE’s complexity is $\mathcal{O}(Ld+P^{2}d)$ where $P\ll L$, compared to $\mathcal{O}(L^{2}d)$ for point-wise models.
+The entropy threshold $\alpha$ provides explicit control: lower $\alpha$ → more boundaries → higher $P$ → more computation but finer granularity.
+Appendix C Anomaly Detection
+Datasets.
+We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD (Su et al., 2019), MSL(Hundman et al., 2018), SMAP (Hundman et al., 2018), SWaT (Mathur and Tippenhauer, 2016), PSM (Abdulaal et al., 2021)) by replacing the forecasting head with an anomaly scoring mechanism.
+Table 4 shows F1 scores compared to recent methods.
+Appendix D Baselines
+Baselines and Experimental Settings.
+Baselines and Experimental Settings.
+We followed the same experimental settings and adopted the same baselines in TSLANet.
+These are TSLANet, GPT4TS, TimesNet, PatchTST, ETSformer (Woo et al., 2022), FEDformer, LightTS (Zhang et al., 2024), DLinear.
+Table 4: Anomaly detection F1 scores (%) on benchmark datasets.
+Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
+SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
+MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
+SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
+SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
+PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
+Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
+Results.
+EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines.
+These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
+D.1 Baseline Details
+See Table 5 for the details about each baseline.
 Table 5: Overview of baseline models (2021–2025) grouped by architecture. We highlight their key design paradigms for deep time series forecasting.
 Model Category Key Features / Innovations
+TimeMosaic (Ding et al., 2025) Transformer patch granularity based on temporal heterogeneity and motif reuse.
 iTransformer (Liu et al., 2024b) Transformer Inverted attention across feature and time dimensions.
@@
 TimesNet (Wu et al., 2023) CNN Temporal 2D variation modeling with convolutional layers.
+TSLANet (Eldele et al., 2024) CNN + Lightweight Lightweight and pretraining module for enhanced feature representation
 TimeMixer (Wang et al., 2024) MLP Multiscale mixing, separation of trend and seasonal signals.
@@
 DLinear (Zeng et al., 2023) Linear/MLP Channel-independent trend/seasonal decomposition.
-TimeKAN (Huang et al., 2025b) Kernel Attn. Network Kernelized attention distinct from Transformer/MLP.
 Time-FFM (Liu et al., 2024a) Foundation Pre-trained foundation model for time series.
@@
 LangTime (Niu et al., 2025) LLM-aligned / Cross-modal Language-guided forecasting with RL optimization.
-CALF (Liu et al., 2025) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
+CALF (Liu et al., 2025b) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
 FilterTS (Wang et al., 2025) Filter-based Static global + dynamic cross-variable frequency filtering.
-A.2 Dataset Details
+D.2 Dataset Details
 Table 6: Summary of datasets used for long-term forecasting evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
@@
 Electricity (ECL) 321 (18317, 2633, 5261) Hourly Electricity Consumption
-Exchange Rate 8 (5120, 665, 1422) Daily Foreign Exchange Market
 This study employs a comprehensive collection of benchmark forecasting datasets to rigorously evaluate model performance across diverse temporal domains and application scenarios.
@@
 The dataset is particularly valuable for optimizing power generation scheduling and distribution network management, representing a critical application domain for time series forecasting in energy systems.
-  • Exchange Rate: This dataset contains daily exchange rates of the US dollar against eight different currencies, spanning from 1990 to 2016.
-It represents financial time series forecasting challenges with inherent volatility and complex market dynamics.
 Table 6 presents the detailed characteristics of all datasets employed in this study.
@@
 Dataset sizes are reported in the format (Training, Validation, Test) to clearly delineate the data partitioning strategy employed across all experiments.
-A.3 Implementation Details
+D.3 Classification Dataset
+For classification we used 10 UEA datasets, following the evaluation and preprocessing protocol of TSLANet (Eldele et al., 2024).
+UEA datasets: We incorporate datasets from the University of East Anglia (UEA) Time Series Classification repository, which is renowned for its rich collection of multivariate time series datasets.
+We were able to preprocess 10 datasets, each offering a multidimensional perspective on time series analysis across various real-world scenarios, such as human activity recognition, sensor data interpretation, and complex system monitoring.
+More details about the UEA datasets can be found in https://www.timeseriesclassification.com/.
+D.4 Anomaly Detection Dataset
+For Anomaly Detection also we followed the evaluation and preprocessing protocol of TSLANet.
+In our study, we assess the performance of our model using five benchmark datasets, each representing a distinct application area, to demonstrate its effectiveness in detecting anomalies in diverse settings:
+  • SMD (Server Machine Dataset): Utilized for server monitoring, the SMD dataset comprises multivariate time series data collected from servers and aims to identify unusual server behaviors that could indicate failures or security issues.
+  • MSL (Mars Science Laboratory): This dataset contains telemetry data from the Mars Science Laboratory rover, focusing on space exploration applications.
+Anomaly detection in this context is crucial for identifying potential issues with spacecraft systems based on their operational data.
+  • SMAP (Soil Moisture Active Passive): Related to earth observations, the SMAP dataset includes soil moisture measurements intended for environmental monitoring.
+Detecting anomalies in soil moisture can provide insights into environmental conditions and potential agricultural impacts.
+  • SWaT (Secure Water Treatment): In the domain of water treatment security, the SWaT dataset consists of data from a water treatment testbed, simulating the operational data of water treatment plants.
+Anomaly detection here is vital for ensuring the safety and security of water treatment processes.
+  • PSM (Pump Sensor Monitoring): Focused on industrial pump sensors, the PSM dataset gathers sensor data from pumps in industrial settings.
+Anomalies in this dataset can indicate equipment malfunctions or the need for maintenance, critical for preventing industrial accidents.
+The detailed characteristics of these datasets is presented in Table 7.
+Table 7: Summary of datasets used for anomaly detection evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
+Dataset Dim Dataset Size (Train, Val, Test) Length Domain Information
+SMD 38 (566724, 141681, 708420) 100 Server Machine
+MSL 55 (44653, 11664, 73729) 100 Spacecraft
+SMAP 25 (108146, 27037, 427617) 100 Spacecraft
+SWaT 51 (396000, 99000, 449919) 100 Infrastructure
+PSM 25 (105984, 26497, 87841) 100 Server Machine
+D.5 Implementation Details
 We summarize the hyperparameter configurations used for EntroPE across all benchmark datasets in Table 9.
@@
 Number of heads (n_headn\_head) 4
-Embedding dimension (n_embdn\_embd) 8
+Embedding dimension (n_embdn\_embd) 16
 Dropout 0.1
@@
 Table 9: Hyperparameter settings for EntroPE across datasets.
-Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity Exchange Rate
-dim 8 8 16 16 16 32 8
-heads 2 2 2 4 2 4 2
-layers 1 2 1 1 2 2 1
-max patch length 24 24 24 24 24 24 16
-batch_size 64 64 32 32 128 32 32
-learning_rate 0.001 0.01 0.01 0.001 0.01 0.01 0.01
-dropout 0.05 0.1 0.1 0.1 0.2 0.1 0.2
-monotonicity 1 1 1 1 1 1 1
-threshold ($\theta$) 3 3.5 3.5 3 3 3.5 3.5
-threshold_add ($\gamma$) 0.25 0.2 0.3 0.25 0.35 0.3 0.3
-train_epochs 20 20 20 20 20 20 20
-patience 7 7 7 7 7 7 7
-A.4 Extended Results
-A.4.1 Full Results (Long-term forecasting). Input length 96.
-We adopted the training criteria established by TimeKAN for fair comparison across all baseline methods.
+Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
+dim 8 8 16 16 16 32
+heads 2 2 2 4 2 4
+layers 1 2 1 1 2 2
+max patch length 24 24 24 24 24 24
+batch_size 64 64 32 32 128 32
+learning_rate 0.001 0.01 0.01 0.001 0.01 0.01
+dropout 0.05 0.1 0.1 0.1 0.2 0.1
+threshold % ($\theta$) 3 3.5 3.5 3 3 3.5
+train_epochs 20 20 20 20 20 20
+Appendix E Extended Experimental Results
+E.1 Time Series Classification
+To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA time series classification datasets.
+We replace the forecasting head with a classification head while keeping the core architecture unchanged.
+Table 10 presents results compared to recent baselines.
+Table 10: Classification accuracy (%) on UEA benchmark datasets. EntroPE achieves competitive or superior performance across diverse classification tasks.
+Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
+EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
+FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
+Handwriting 57.4 57.88 3.76 26.18 48.47 26.24 26.00 22.47 47.76
+Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
+JapaneseVowels 99.0 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
+PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
+SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
+SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
+SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
+UWaveGestureLibrary 94.4 91.25 84.38 83.13 94.40 81.56 80.00 81.88 86.25
+Average 77.13 76.03 64.37 69.36 72.90 70.88 71.14 69.51 73.53
+EntroPE achieves an average accuracy of 77.13%, outperforming TSLANet (76.03%), TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines.
+This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
+E.2 Anomaly Detection
+We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD, MSL, SMAP, SWaT, PSM) by replacing the forecasting head with an anomaly scoring mechanism.
+Table 11 shows F1 scores compared to recent methods.
+Table 11: Anomaly detection F1 scores (%) on benchmark datasets. EntroPE demonstrates strong performance across diverse anomaly detection scenarios.
+Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
+SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
+MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
+SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
+SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
+PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
+Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
+EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines.
+These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
+E.3 Comparison with Large-Scale Foundation Models
+We compare EntroPE with recent large-scale time series foundation models including MOIRAI (Small, Base, Large), Sundial (Small, Base, Large), and TimeMoE (Base, Large, XL).
+Table 12 presents results on ETT and related datasets.
+Table 12: Comparison with large-scale foundation models (MSE / MAE). EntroPE achieves competitive performance with significantly fewer parameters.
+Dataset EntroPE MOIRAI_S MOIRAI_B MOIRAI_L Sundial_S Sundial_B Sundial_L TimeMoE_B TimeMoE_L TimeMoE_XL
+ETTm1 0.378/0.391 0.400/0.424 0.434/0.439 0.510/0.468 0.354/0.388 0.336/0.377 0.331/0.369 0.394/0.415 0.376/0.405 0.341/0.382
+ETTm2 0.286/0.335 0.311/0.348 0.279/0.330 0.282/0.330 0.265/0.324 0.258/0.320 0.254/0.315 0.317/0.365 0.316/0.361 0.288/0.344
+ETTh1 0.416/0.425 0.400/0.423 0.434/0.439 0.510/0.468 0.390/0.418 0.411/0.434 0.395/0.420 0.400/0.424 0.394/0.409 0.412/0.426
+ETTh2 0.366/0.387 0.348/0.374 0.354/0.376 0.359/0.371 0.340/0.387 0.333/0.387 0.334/0.387 0.366/0.404 0.405/0.415 0.371/0.399
+Electricity 0.182/0.271 0.238/0.303 0.219/0.283 0.229/0.281 0.169/0.265 0.169/0.265 0.166/0.262
+Weather 0.242/0.273 0.233/0.271 0.234/0.270 0.238/0.275 0.233/0.271 0.234/0.270 0.238/0.275 0.265/0.297 0.270/0.300 0.256/0.288
+Despite having orders of magnitude fewer parameters (EntroPE: $\sim$0.1-0.2M vs.
+foundation models: 10M-500M+), EntroPE achieves competitive or superior performance on several datasets, demonstrating the effectiveness of entropy-guided dynamic patching as a lightweight alternative to massive pre-trained models.
+E.4 Comprehensive Efficiency Comparison
+Table 13 presents a detailed comparison of parameters and MACs (Multiply-Accumulate Operations) across multiple datasets for EntroPE and recent baselines.
+Table 13: Comprehensive efficiency analysis: Parameters and MACs across datasets. EntroPE achieves competitive accuracy with significantly lower computational cost.
+Model ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
+Params MACs Params MACs Params MACs Params MACs Params MACs Params MACs
+TimeMixer 75.50K 20.37M 75.50K 20.37M 75.50K 20.37M 77.77K 24.18M 104.43K 82.62M 106.83K 1.26G
+iTransformer 841.57K 77.46M 224.22K 19.86M 224.22K 19.86M 224.22K 19.86M 4.83M 1.16G 4.83M 16.29G
+PatchTST 3.75M 5.90G 10.06M 17.66G 3.75M 5.90G 10.06M 17.66G 6.90M 35.30G 6.90M 539.38G
+TimesNet 605.48K 18.13G 1.19M 36.28G 4.71M 144G 1.19M 36.28G 1.19M 36.28G 150.30M 4.61T
+MICN 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 111.03M 295.07M 6.64M 19.5G
+FiLM 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 8.46G 12.58M 8.46G
+FEDFormer 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.45M 25.23G 24.29M 30.89G
+AutoFormer 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.61M 22.98G 12.14M 28.75G
+EntroPE 95K 9M 418K 53M 192K 20M 95K 10M 195K 61M 199K 1G
+[…]
+E.5 EntroPE as Plug-in Replacement for PatchTST
+To demonstrate the modularity and effectiveness of entropy-guided patching, we integrate EntroPE’s dynamic patcher and Fusion Decoder into PatchTST while keeping other components unchanged.
+Table 14 shows the performance improvement.
+Table 14: EntroPE as plug-in replacement for PatchTST’s static patcher. Relative improvement (%) shows consistent gains across datasets.
+Model ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
+MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+PatchTST 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
+EntroPE 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
+$\Delta$% Improvement 19% 12% 6% 6% 7% 4% 1% 0% 9% 4% 16% 15%
+Replacing PatchTST’s static patcher with EntroPE’s entropy-guided dynamic patcher yields approximately 8-10% average MSE improvement across datasets, reduces parameters from $\sim$4M to 50K-1M range, and decreases training time.
+This demonstrates that EntroPE can serve as a drop-in replacement for static patching mechanisms in existing architectures.
+E.6 Training Efficiency Analysis
+We measure end-to-end training time per epoch to assess the computational overhead of dynamic patching. Table 15 shows that entropy-guided patching adds minimal overhead.
+Table 15: Training efficiency comparison: Epoch duration (seconds) for static vs. dynamic patching.
+Metric/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
+Epoch (dynamic) 3.2 3.5 17.2 20.8 54.5 181
+Epoch (static) 3.1 3.4 17.0 20.5 54.1 180
+Patching time 0.1 0.1 0.2 0.3 0.4 1
+Overhead % 3% 3% 1% 1% 1% 1%
+Dynamic patching increases epoch time by only 0.1-1.0 seconds (1-3% overhead), demonstrating that the entropy computation and boundary detection are computationally negligible compared to the transformer’s forward-backward passes.
+E.6.1 Full Results (Long-term forecasting). Input length 96.
+We adopted the training criteria established by TimeKAN and TimesNet for fair comparison across all baseline methods.
 The results for AutoFormer, FEDformer, TimesNet, PatchTST, Time-FFM, iTransformer, and TimeMixer were directly extracted from the TimeKAN paper to ensure consistency in experimental conditions.
-For HDMixer, TimeBase, CALF, and FilterTS, we executed the original implementations following identical training criteria.
+For HDMixer, TimeBase, CALF, TSLANet, TimeMosaic and FilterTS we executed the original implementations following identical training criteria.
 Specifically for FilterTS, we addressed the challenge of multiple configuration options by running experiments across all different settings proposed by the authors for each forecasting horizon.
-We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section A.4.2).
+We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section E.6.2).
 LangTime baseline results were obtained directly by running the author’s implementation to maintain consistency with the their reported performance.
@@
 Table 16: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model. Look-back window $L$ set to 96 for all cases, forecast window $T=\{96,192,336,720\}$
-Dataset T EntroPE TimeKAN HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
+Dataset T EntroPE TimeMosaic HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
 MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.317 0.354 0.327 0.365 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
-192 0.360 0.379 0.354 0.380 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
-336 0.388 0.399 0.386 0.403 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
-720 0.446 0.433 0.452 0.442 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
-Avg. 0.378 0.391 0.380 0.398 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
-ETTm2 96 0.180 0.265 0.175 0.258 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
-192 0.241 0.311 0.241 0.301 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
-336 0.313 0.352 0.308 0.349 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
-720 0.411 0.413 0.417 0.414 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
-Avg. 0.286 0.335 0.285 0.331 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
-ETTh1 96 0.375 0.399 0.367 0.395 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
-192 0.423 0.425 0.416 0.424 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
-336 0.429 0.432 0.446 0.436 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
-720 0.439 0.454 0.441 0.453 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
+ETTm1 96 0.317 0.354 0.312 0.351 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
+192 0.360 0.379 0.370 0.379 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
+336 0.388 0.399 0.386 0.399 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
+720 0.446 0.433 0.469 0.440 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
+Avg. 0.378 0.391 0.388 0.392 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
+ETTm2 96 0.180 0.265 0.177 0.254 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
+192 0.241 0.311 0.240 0.295 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
+336 0.313 0.352 0.298 0.336 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
+720 0.411 0.413 0.397 0.391 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
+Avg. 0.286 0.335 0.278 0.319 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
+ETTh1 96 0.375 0.399 0.369 0.397 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
+192 0.423 0.425 0.419 0.439 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
+336 0.429 0.432 0.454 0.451 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
+720 0.439 0.454 0.466 0.479 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
 Avg. 0.416 0.425 0.418 0.427 0.448 0.437 0.463 0.429 0.441 0.435 0.440 0.432 0.437 0.425 0.459 0.444 0.442 0.434
-ETTh2 96 0.281 0.336 0.290 0.339 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
-192 0.371 0.393 0.386 0.399 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
-336 0.392 0.394 0.441 0.447 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
-720 0.421 0.427 0.450 0.458 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
-Avg. 0.366 0.387 0.391 0.410 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
-Weather 96 0.164 0.211 0.163 0.209 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
-192 0.210 0.252 0.208 0.251 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
-336 0.256 0.290 0.263 0.290 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
-720 0.339 0.342 0.341 0.341 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
-Avg. 0.242 0.273 0.244 0.273 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
-Electricity 96 0.163 0.252 0.174 0.266 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
-192 0.177 0.268 0.182 0.273 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
-336 0.194 0.284 0.197 0.286 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
-720 0.235 0.321 0.236 0.320 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
-Avg. 0.182 0.271 0.197 0.286 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299
-Exchange 96 0.083 0.140 0.084 0.201 0.089 0.201 0.087 0.205 0.081 0.201
-192 0.177 0.235 0.173 0.295 0.175 0.298 0.179 0.300 0.168 0.293
-336 0.299 0.371 0.316 0.407 0.329 0.409 0.333 0.417 0.299 0.396
-720 0.761 0.869 0.827 0.685 0.852 0.690 0.912 0.719 0.805 0.674
-Avg. 0.331 0.386 0.355 0.400 0.361 0.400 0.378 0.410 0.338 0.391
+ETTh2 96 0.281 0.336 0.292 0.341 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
+192 0.371 0.393 0.367 0.391 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
+336 0.392 0.394 0.417 0.427 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
+720 0.421 0.427 0.398 0.436 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
+Avg. 0.366 0.387 0.369 0.399 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
+Weather 96 0.164 0.211 0.166 0.201 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
+192 0.210 0.252 0.219 0.258 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
+336 0.256 0.290 0.273 0.296 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
+720 0.339 0.342 0.354 0.351 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
+Avg. 0.242 0.273 0.253 0.277 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
+Electricity 96 0.163 0.252 0.161 0.261 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
+192 0.177 0.268 0.176 0.274 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
+336 0.194 0.284 0.189 0.283 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
+720 0.235 0.321 0.225 0.304 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
+Avg. 0.182 0.271 0.188 0.281 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299
 Table 17: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model.
@@
 (2025b).
-Dataset T EntroPE iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.317 0.354 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
-192 0.360 0.379 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
-336 0.388 0.399 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
-720 0.446 0.433 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
-Avg. 0.378 0.391 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
-ETTm2 96 0.180 0.265 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
-192 0.241 0.311 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
-336 0.313 0.352 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
-720 0.411 0.413 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
-Avg. 0.286 0.335 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
-ETTh1 96 0.375 0.399 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
-192 0.423 0.425 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
-336 0.429 0.432 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
-720 0.439 0.454 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
-Avg. 0.416 0.425 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
-ETTh2 96 0.281 0.336 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
-192 0.371 0.393 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
-336 0.392 0.394 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
-720 0.421 0.427 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
-Avg. 0.366 0.387 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
-Weather 96 0.164 0.211 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
-192 0.210 0.252 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
-336 0.256 0.290 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
-720 0.339 0.342 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
-Avg. 0.242 0.273 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
-Electricity 96 0.163 0.252 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
-192 0.177 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
-336 0.194 0.284 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
-720 0.235 0.321 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
-Avg. 0.182 0.271 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
-Exchange 96 0.083 0.140 0.086 0.206 0.088 0.205 0.107 0.234 0.088 0.218 0.148 0.278 0.197 0.323
-192 0.177 0.235 0.177 0.299 0.176 0.299 0.226 0.344 0.176 0.315 0.271 0.315 0.300 0.369
-336 0.299 0.371 0.331 0.417 0.301 0.397 0.367 0.448 0.313 0.427 0.460 0.427 0.509 0.524
-720 0.761 0.869 0.847 0.691 0.901 0.714 0.964 0.746 0.839 0.695 1.195 0.695 1.447 0.941
-Avg. 0.331 0.386 0.360 0.403 0.367 0.404 0.416 0.443 0.354 0.414 0.519 0.429 0.613 0.539
-A.4.2 FilterTS Performance Comparison Across Configurations
+Dataset T EntroPE TSLANet iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
+MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm1 96 0.317 0.354 0.321 0.362 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
+192 0.360 0.379 0.361 0.383 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
+336 0.388 0.399 0.382 0.404 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
+720 0.446 0.433 0.446 0.438 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
+Avg. 0.378 0.391 0.378 0.397 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
+ETTm2 96 0.180 0.265 0.179 0.261 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
+192 0.241 0.311 0.243 0.303 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
+336 0.313 0.352 0.308 0.345 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
+720 0.411 0.413 0.403 0.400 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
+Avg. 0.286 0.335 0.283 0.327 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
+ETTh1 96 0.375 0.399 0.387 0.405 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
+192 0.423 0.425 0.448 0.436 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
+336 0.429 0.432 0.451 0.437 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
+720 0.439 0.454 0.505 0.485 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
+Avg. 0.416 0.425 0.448 0.441 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
+ETTh2 96 0.281 0.336 0.299 0.354 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
+192 0.371 0.393 0.382 0.399 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
+336 0.392 0.394 0.379 0.396 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
+720 0.421 0.427 0.429 0.445 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
+Avg. 0.366 0.387 0.372 0.399 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
+Weather 96 0.164 0.211 0.177 0.217 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
+192 0.210 0.252 0.224 0.257 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
+336 0.256 0.290 0.278 0.297 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
+720 0.339 0.342 0.355 0.347 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
+Avg. 0.242 0.273 0.259 0.280 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
+Electricity 96 0.163 0.252 0.175 0.260 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
+192 0.177 0.268 0.182 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
+336 0.194 0.284 0.199 0.285 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
+720 0.235 0.321 0.240 0.317 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
+Avg. 0.182 0.271 0.199 0.283 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
+E.6.2 FilterTS Performance Comparison Across Configurations
 The original FilterTS results (18) were reported using different hyperparameter settings for different forecast lengths, which complicates direct comparison with other baselines.
@@
 Avg. 0.184 0.275 0.184 0.275
-Exchange Rate 96 96 0.085 0.201 0.083 0.200 0.084 0.201 0.082 0.198
-96 192 0.180 0.300 0.171 0.294 0.173 0.295 0.172 0.294
-96 336 0.342 0.420 0.323 0.411 0.316 0.407 0.322 0.410
-96 720 0.878 0.699 0.916 0.713 0.827 0.685 0.828 0.686
-Avg. 0.371 0.405 0.373 0.405 0.350 0.397 0.351 0.397
-A.5 Entropy Qualitative Analysis
+Appendix F Entropy Qualitative Analysis
 Entropy Visualization: Figure 6 demonstrates how entropy-driven boundaries align with temporal transitions, creating patches that preserve intra-patch coherence while capturing natural temporal structure.
@@
 Figure 6: Comprehensive entropy analysis for dynamic patching on ETT and Electricity data. Subfigures show how entropy-driven patches align with temporal transitions.
-A.6 Time Series Quantization
+F.1 Time Series Quantization
 Real-valued time series data cannot be directly processed by entropy-based boundary detection methods that rely on discrete probability distributions.
@@
 Unlike traditional approaches that rely on dequantization for prediction, our method uses quantization purely for entropy-based boundary detection while employing linear projection on continuous representations for forecasting.
-A.7 Entropy Calculation
+Automatic Determination of Quantization Range.
+We avoid manually choosing a fixed tokenization range (e.g., [8,8][-8,8]) and instead determine the quantization interval automatically from the training data.
+This ensures that the discretization step is fully data-driven and does not require hand-tuning for each dataset.
+$\gamma_{\text{rel}}=Q_{\alpha}\left({\Delta H(x_{t})}_{t=2}^{T}\right)$ (39)
+First, all continuous features are z-score normalized using the mean and standard deviation computed from the training split only.
+Let $X_{\text{train}}$ denote the collection of all normalized values that are passed to the tokenizer.
+We compute two empirical quantiles:
+$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\qquad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}}),$
+where $\epsilon$ controls the desired coverage (we use $\epsilon=0.005$, corresponding to $99.5\%$ of the training distribution). We then define a symmetric quantization radius
+$R=\max\left(|q_{\text{low}}|,\;|q_{\text{high}}|\right),$
+and construct the discretization interval as [R,R][-R,R].
+This interval is uniformly divided into $V$ bins (we use $V=256$ unless otherwise stated).
+Values falling outside [R,R][-R,R] are clipped to the nearest boundary bin.
+This procedure allows the tokenizer to adapt automatically to the empirical scale and variability of the dataset, providing a robust discretization scheme without manual hyperparameter selection.
+Appendix G Utility & Loss Functions
+G.1 Entropy Calculation
 We begin with the standard Shannon entropy for a discrete random variable $X$ with possible values $\{x_{1},x_{2},\ldots,x_{n}\}$:
@@
 This conditional entropy quantifies the uncertainty in predicting the next value given the historical context, with higher values indicating less predictable (more informative) regions of the time series.
-A.8 Cross-Entropy Loss
+G.2 Cross-Entropy Loss
 $L_{CE}=-\sum_{i=1}^{N-1}\log P(t_{i+1}|t_{1},\ldots,t_{i};\theta)$ (43)
 where $\theta$ represents the model parameters.
-A.9 Mean Squared Error (MSE)
+G.3 Mean Squared Error (MSE)
 $L_{MSE}=\frac{1}{N}\sum_{i=1}^{N}({y_{i}-\hat{y_{i}}})^{2}$
-A.10 EntroPE Model Algorithms
-Algorithm 1 EntroPE Forward Pass
-1:Raw time series $X\in\mathbb{R}^{B\times C\times L}$, forecast horizon $T$
-2:Predictions $\hat{Y}\in\mathbb{R}^{B\times C\times T}$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-14:return $\hat{Y}$
-Algorithm 2 Entropy-Based Dynamic Patcher (EDP)
-1:Tokenized sequence $X^{tok}\in\mathbb{R}^{(B\cdot C)\times L}$, pre-trained entropy model $\theta$
-2:Patch boundaries $\mathcal{B}$, patch mask $M\in\mathbb{R}^{(B\cdot C)\times L}$
-3:Initialize $\mathcal{B}\leftarrow\emptyset$, $M\leftarrow\mathbf{0}$
-4:Phase 1: Entropy Calculation
-5:for $t=1$ to L1L-1 do
-6:  $\text{context}\leftarrow X^{tok}[:,1:t]$
-7:  $\text{logits}\leftarrow\text{entropy\_model}_{\theta}(\text{context})$
-8:  $p_{\theta}(x_{t+1}|x_{\leq t})\leftarrow\text{softmax}(\text{logits})$
-9:  $H(x_{t})\leftarrow-\sum_{v\in\mathcal{V}}p_{\theta}(v)\log p_{\theta}(v)$
-10:end for
-11:Phase 2: Boundary Detection
-12:for $t=2$ to L1L-1 do
-13:  if $H(x_{t})>\theta$ and $H(x_{t})-H(x_{t-1})>\gamma$ then
-14:   $\mathcal{B}\leftarrow\mathcal{B}\cup\{t\}$;  $M[:,t]\leftarrow 1$
-15:  end if
-16:end for
-17:Add first and last positions as boundaries
-[…]
-19:M[:,1]1M[$\leftarrow 1$;  $M[:,L]\leftarrow 1$
-20:return $\mathcal{B},M$
-Algorithm 3 Adaptive Patch Encoder (APE)
-1:Continuous sequence $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$, patch boundaries $\mathcal{B}$, vocab size $V$
-2:Patch embeddings $P^{in}$, hidden states $H^{enc}$
-[…]
-[…]
-5:Extract variable-length patches
-[…]
-7:for $i=1$ to $|\mathcal{B}|-1$ do
-8:  $\text{patch}_{i}\leftarrow X^{ci}[:,\mathcal{B}[i]:\mathcal{B}[i+1]]$
-9:  Append $\text{patch}_{i}$ to patches
-10:end for
-11:Initial patch embedding via max pooling
-12:for each patch $p_{j}$ do
-13:  $\text{emb}\leftarrow E(\text{tokenize}(p_{j}))$
-14:  $P_{0}[:,j,:]\leftarrow\text{MaxPool}(\text{emb})$
-15:end for
-16:Cross-attention refinement (NN layers)
-[…]
-18:for $n=1$ to NN do
-19:  $h_{n}\leftarrow\text{TransformerEncoder}_{n}(h_{\text{curr}})$
-20:  $Q\leftarrow W_{q}(P_{\text{curr}})$
-21:  $K\leftarrow W_{k}(h_{n})$,   $V\leftarrow W_{v}(h_{n})$
-22:  $\text{out}\leftarrow\text{softmax}\!\left(\tfrac{QK^{T}}{\sqrt{d_{k}}}\right)V$
-23:  $P_{n}\leftarrow P_{\text{curr}}+W_{o}(\text{out})$
-24:  Update $h_{\text{curr}},P_{\text{curr}}$
-25:end for
-26:return $P^{in}\leftarrow P_{N}$,   $H^{enc}\leftarrow h_{N}$
-A.11 Use of Large Language Models
+Use of Large Language Models
 Parts of the text in this paper were refined with the assistance of a large language model (ChatGPT, GPT-5), used exclusively for language polishing and improving clarity of exposition.